### v1

# On the 2-torsion in class groups of number fields

Abstract.
In $2020$, Bhargava, Shankar, Taniguchi, Thorne, Tsimerman, and Zhao proved that for a finite extension $K/\mathbb{Q}$ of degree $n\geq 5$, the size of the 22-torsion class group is bounded by $h_{2}(K)=O_{n,\varepsilon}(D_{K}^{\frac{1}{2}-\frac{1}{2n}+\varepsilon})$, where $D_{K}$ is the absolute discriminant of $K$. In the present paper, we improve their bound by proving that $h_{2}(K)=O_{n,\varepsilon}(D_{K}^{\frac{1}{2}-\frac{1}{2n}-\delta_{K}+\varepsilon})$, for a constant $\delta_{K}\geq\frac{1}{28n}-\frac{3}{28n(n-1)}$.
1. Introduction
For a given number field $K/\mathbb{Q}$ of degree $n$ and for any $\ell$ prime, a classical problem in algebraic number theory is to bound the size of $h_{\ell}(K)$, the $\ell$-torsion subgroup of the class group $h(K)$. Since $h_{\ell}(K)\leq\#h(K)$, from the Brauer-Siegel Theorem (see [Lan94, Chapter XVI]) one deduces that
$h_{\ell}(K)=O_{n,\varepsilon}(D_{K}^{1/2}(\log D_{K})^{n-1}),$
where $D_{K}$ denotes the absolute value of the discriminant of $K$ over $\mathbb{Q}$. Recently, Lemke Oliver and Zaman ([LZ25, Corollary 11]) showed that for every $\ell$, one has $h_{\ell}(K)=o(D_{K}^{\frac{1}{2}})$ (and where the saving is an explicit power of $\log(D_{K})$). On the other hand, it is conjectured that
$h_{\ell}(K)=O_{n,\varepsilon}(D_{K}^{\varepsilon}),$ (1)
for any $\varepsilon>0$. This conjecture remains widely open; the only case currently known is $(n,\ell)=(2,2)$, which follows from Gauss’ genus theory. In fact, even establishing the existence of a constant $\delta_{n,\ell}>0$ such that for any number field $K$ of degree $n$, we have
$h_{\ell}(K)\ll D_{K}^{\frac{1}{2}-\delta_{n,\ell}+\varepsilon}$
It is already a challenging problem. In this paper, we focus on the case $\ell=2$. It was shown in [BSTTZ20, Theorem 1.1] that one may take
$\delta_{n,2}=\begin{cases}0.2874\ldots&\text{if }n=3,4,\\ \frac{1}{2n}&\text{if }n\geq 5.\end{cases}$
Our main result is the following:
Theorem $1.$
Let $K/\mathbb{Q}$ be a number field of degree $n\geq 5$. Then, there exists $\delta_{K}>0$ such that for any $\varepsilon>0$, one has
$h_{2}(K)=O_{n,\varepsilon}(D_{K}^{\frac{1}{2}-\frac{1}{2n}-\delta_{K}+\varepsilon}).$
Moreover we have that $\delta_{K}\geq\frac{1}{28n}-\frac{3}{28n(n-1)}$.
2. Outline of the proof
We start by fixing some notation: let $\mathcal{O}_{K}$ be the ring of integers of $K$. It is a well-known fact that $\mathcal{O}_{K}$ is a $\mathbb{Z}$-lattice of rank $n$. Moreover, we have $\text{Disc}(\mathcal{O}_{K})=D_{K}^{\frac{1}{2}}$. For every $\alpha\in\mathcal{O}_{K}$, consider the norm:
$\|\alpha\|=\sum_{\sigma\in\operatorname{Emb}(K/\mathbb{Q})}|\sigma(\alpha)|,$
where $\operatorname{Emb}(K/\mathbb{Q})$ denotes the set of embeddings of $K$ into $\mathbb{C}$ fixing $\mathbb{Q}$. Since $\mathcal{O}_{K}$ is a lattice, we can find a basis $\{\alpha_{1,K}=1,\alpha_{2,K},\ldots,\alpha_{n,K}\}$ which is Minkowski-reduced, i.e. such that $1\leq\|\alpha_{2,K}\|\leq\cdots\leq\|\alpha_{n,K}\|$, and for each $i\leq n-1$, $\|\alpha_{i,K}\|\ll\|v\|$ for every $v\notin\text{Span}\{1,\alpha_{2,K},\cdots,\alpha_{i-1,K}\}$. Observe that this implies that
$\|\alpha_{2,K}\|\cdots\|\alpha_{n,K}\|\asymp D_{K}^{\frac{1}{2}}.$ (2)
Finally, for each i=2,,ni=2,...,n, we denote by $\delta_{i,K}$ the real positive number such that $\|\alpha_{i,K}\|=D_{K}^{\delta_{i,K}}$. At this point, it is useful to record the following
Proposition 2.1.
Let $K/\mathbb{Q}$ be a finite extension of degree $n$. Then
  • $i)$
    $\|\alpha_{n,K}\|=O_{n}(D_{K}^{\frac{1}{n}})$. In particular we have $0\leq\delta_{i,K}\leq\frac{1}{n}+o(1)$, for each $2\leq i\leq n$;
  • ii)ii)
    $\|\alpha_{2,K}\|=O_{n}(D_{K}^{\frac{1}{2(n-1)}})$. In particular we have $0\leq\delta_{2,K}\leq\frac{1}{2(n-1)}+o(1)$.
Proof.
We start with $(i)$. The bound $\|\alpha_{n,K}\|=O_{n}(D_{K}^{\frac{1}{n}})$ is proved in [BSTTZ20, Theorem $1.6$]. Since the basis $\{1,\alpha_{2,K},\ldots,\alpha_{n,K}\}$ is Minkowski-reduced, we have $\|\alpha_{i,K}\|\leq\|\alpha_{n,K}\|=O(D_{K}^{\frac{1}{n}})$ for every $i\in\{2,...,n\}$. Hence, $\delta_{i,K}\leq\frac{1}{n}+o(1)$ for every $i\in\{2,...,n\}$, as claimed.
[…]
$D_{K}^{\frac{1}{2}}\gg_{n}\|\alpha_{2,K}\|\cdots\|\alpha_{n,K}\|\geq\|\alpha_{2,K}\|^{n-1}.$
Our starting point is the same as [BSTTZ20, Theorem $1.1$]: The size of the 22-torsion of a number field extension is bounded by
$h_{2}(K)\ll\#\left\{(y,z_{1},...,z_{n})\in\mathbb{Z}^{n+1}:\begin{matrix}|z_{i}|\ll\frac{D_{K}^{\frac{1}{n}}}{\|\alpha_{i,K}\|},\quad\text{for }i=1,...,n\\ y^{2}=N_{K}(z_{1}+z_{2}\alpha_{2,K}+\cdots+z_{n}\alpha_{n,K})\end{matrix}\right\}.$
In [BSTTZ20], the bound for $h_{2}(K)$ (when $n\geq 5$) is achieved by fixing a choice of $z_{2},...,z_{n}$ and by using the Bombieri-Pila bound ([BP89]) to count integral solutions $(y,x_{1})$ of the curve,
$Y^{2}=N_{K}(X_{1}+z_{2}\alpha_{2,K}+\cdots+z_{n}\alpha_{n,K}),$
in the box $D_{K}^{\frac{1}{2}},D_{K}^{\frac{1}{2}}]\times[-D_{K}^{\frac{1}{n}},D_{K}^{\frac{1}{n}}]$. Similarly to Salberger’s approach ([Sal22]), we shall instead fix only $z_{3},...,z_{n}$ and count integral solutions $(y,x_{1},x_{2})$ on the surface
$Y^{2}=N_{K}(X_{1}+X_{2}\alpha_{2,K}+z_{3}\alpha_{3,K}\cdots+z_{n}\alpha_{n,K}),$
in the box
$D_{K}^{\frac{1}{2}},D_{K}^{\frac{1}{2}}]\times[-D_{K}^{\frac{1}{n}},D_{K}^{\frac{1}{n}}]\times[-D_{K}^{\frac{1}{n}-\delta_{2,K}},D_{K}^{\frac{1}{n}-\delta_{2,K}}].$ (3)
[…]
2.1. The case $\delta_{2,K}\leq\frac{1}{2n}$.
In this case, it will be enough to deploy the classical square sieve as introduced in [HB84], getting
Theorem $2.$
Let $n\geq 5$ and $\varepsilon>0$. For any number field $K$ of degree $n$, we have
$h_{2}(K)\ll_{n,\varepsilon}D_{K}^{\frac{1}{2}-\frac{1}{2n}-\left(\frac{1}{6n}-\frac{\delta_{2,K}}{3}\right)+\varepsilon}$
That is, when $0\leq\delta_{2,K}\leq\frac{1}{2n}$, Theorem 1.1 holds for $\delta_{K}=\frac{1}{6n}-\frac{\delta_{2,K}}{3}$. Moreover, since $\frac{1}{6n}-\frac{\delta_{2,K}}{3}<0$ when $\delta_{2,K}\geq\frac{1}{2n}$, it follows the bound in Theorem 2.2 is actually worst then the one in [BSTTZ20, Theorem $1.1$] in the range $\frac{1}{2n}\leq\delta_{2,K}\leq\frac{1}{2(n-1)}+o(1)$; hence, Theorem 2.2 alone would not suffice to prove Theorem $1.$
$2.$ The case $\frac{1}{2n}\leq\delta_{2,K}\leq\frac{1}{2(n-1)}+o(1)$.
As mentioned earlier, this case will be the most complicated due to the fact that the range of the variable $X_{2}$ is considerably smaller than the one of $X_{1}$. To address this, we combine the square sieve with the $q$-analogue of van der Corput (as in [Pie06]), proving
Theorem 2.3.
Let $n\geq 5$ be an integer and $\varepsilon>0$. For any number field $K$ of degree $n$, we have
$h_{2}(K)\ll_{n,\varepsilon}D_{K}^{\frac{1}{2}-\frac{1}{2n}-\left(\frac{1}{7n}-\frac{3\delta_{2,K}}{14}\right)+\varepsilon},$
provided that $\frac{1}{4n}\leq\delta_{2,K}$.
That is, when $\frac{1}{4n}\leq\delta_{2,K}\leq\frac{1}{2(n-1)}+o(1)$, Theorem 1.1 holds for $\delta_{K}=\frac{1}{7n}-\frac{3\delta_{2,K}}{14}$. Thus, by defining
$\delta_{K}=\max\left\{\frac{1}{6n}-\frac{\delta_{2,K}}{3},\frac{1}{7n}-\frac{3\delta_{2,K}}{14}\right\}\geq\frac{1}{28n}-\frac{3}{28n(n-1)},$
one finally proves Theorem $1.$
The main ingredient in proving Theorem 2.2 is the following stratification result for exponential sums
Proposition 2.4.
Let $F(X,Y)\in\mathbb{F}_{p}[X,Y]$ a square free polynomial such that
$F[X,Y]=\prod_{i=1}^{n}(X+a_{i}Y+b_{i})\text{ over }\overline{\mathbb{F}}_{p},$
and assume that the following conditions hold:
  • $i)$
    There exists $i\in\{2,...,n\}$ such that $a_{1}\neq a_{i}$.
  • ii)ii)
    $\frac{a_{i}-a_{1}}{a_{j}-a_{1}}\cdot(b_{j}-b_{1})\neq b_{i}-b_{1}$ for any $i,j$ such that $a_{j}\neq a_{1}$ and $i\neq j$.
  • $iii)$
    For every $i,j$ with $i\neq j$, $b_{i}\neq b_{j}$.
Then there exists $A_{s}\subset\{(r,t)\in\mathbb{F}_{p}\times\mathbb{F}_{p}\}$ such that
  • $i)$
    for any $(r,t)\in(\mathbb{F}_{p}\times\mathbb{F}_{p})\setminus A_{s}$, one has
    $\left|\sum_{(x,y,z)\in\mathbb{F}_{p}^{3}}\left(\frac{F(x,y)}{p}\right)\left(\frac{F(z,y+r)}{p}\right)e_{p}(s(x-z)+ty)\right|\ll_{n}p^{3/2}(s,t,p)^{1/2}.$
  • ii)ii)
    $A_{s}=O_{n}(1)$.
When $(s,p)=1$, Proposition 2.4, implies that the exponential sum
$\sum_{(x,y,z)\in\mathbb{F}_{p}^{3}}\left(\frac{F(x,y)}{p}\right)\left(\frac{F(z,y+r)}{p}\right)e_{p}(s(x-z)+ty)$ (4)
exhibits square root cancellation for all $(r,t)$ which are not contained in a set of codimension 0 in $\mathbb{F}_{p}^{2}$. Then Proposition 2.4 is the analogue of showing that the AA-number associated to the exponential sum 4, in the sense of [Kat89], is not zero. We remark that, in general, determining whether the AA-number associated with an exponential sum does not vanishes is a difficult problem (see [Kat94], [FK01]).
3. Previous results
[…]
$h_{\ell}(K)\ll_{n,\varepsilon}D_{K}^{\frac{1}{2}-\delta_{n,\ell}+\varepsilon}.$
We can summarize all the cases in which some power-saving bound has been achieved in the following table:
[…]
Gauss 1/2 2 2
[Pie05, Pie06] $1/56$ 2 3
[HV06] $0.0582\ldots$ 2 3
[EV07] $1/6$ 2, 3 3
[CK25] $0.1806$ 2 3
[HB24] $1/(2\ell)$ 2, 3 prime $p\mid D_{K}\Rightarrow p\leq D_{K}^{\delta_{\ell}}$
[HB24] $1/(4\ell)$ 3 prime pure cubic extension
[BSTTZ20] $0.2215\ldots$ 3, 4 2
[EV07] $1/168$ 4 3 $\mathrm{Gal}(K/\mathbb{Q})\cong A_{4}$ or $S_{4}$
[EV07] $\delta$ 4 3 $\mathrm{Gal}(K/\mathbb{Q})\cong C_{4}$, $V_{4}$, or $D_{4}$
[BSTTZ20] $1/(2n)$ $\geq 5$ 2
[Sal22] $\frac{2-n\delta_{2,K}+2\sqrt{1-2n\delta_{2,K}}}{n}$ $\geq 5$ 2 $\delta_{2,K}\leq 1/(2n)$
[KW22] 1/2 $p^{r}$ $p^{s}$ $\mathrm{Gal}(K/\mathbb{Q})\cong G$, any $p$-group $G$
[Wan21, Wan20] $\delta_{G,\ell}$ powerful >2>2 $\mathrm{Gal}(K/\mathbb{Q})\cong G$, many nilp. grps $G$
We remark that Theorem 2.2 improves upon [Sal22] for all number field $K$ with $0\leq\delta_{2,K}\leq\frac{1}{2n}$. Significant efforts have been made to achieve unconditional improvements on the trivial bound on average across families of number fields: [HP17], [EPW17], [Wid18], [FW18], [PTBW20], [An20], [FW21], [TZ23], [LOTZ24], [KT24], [LOWW25], and [CK25].
3.1. Some application of Therem 2.3.
We conclude the intrudiction by presenting some consequences of Theorem 2.3.
Corollary 3.1.
Let $\mathbb{Q}\subset F$ be a finite extension of $\mathbb{Q}$. Then for every $\mathbb{Q}\subset F\subset K$ extension of degree $n$, we have
$h_{2}(K)\ll_{D_{F},n,\varepsilon}D_{K}^{\frac{1}{2}-\frac{2}{3n}+\varepsilon}.$
Proof.
Since $F\subset K$, one has that $\|\alpha_{2,K}\|\leq\|\alpha_{2,F}\|$. Hence, Theorem 2.3 implies
$h_{2}(K)\ll_{n,\varepsilon}D_{K}^{\frac{1}{2}-\frac{1}{2n}-\left(\frac{1}{6n}-\frac{\delta_{2,K}}{3}\right)+\varepsilon}=D_{K}^{\frac{1}{2}-\frac{2}{3n}+\varepsilon}\|\alpha_{2,K}\|^{\frac{1}{3}}\leq D_{K}^{\frac{1}{2}-\frac{2}{3n}+\varepsilon}\|\alpha_{2,F}\|^{\frac{1}{3}}\ll_{D_{F}}D_{K}^{\frac{1}{2}-\frac{2}{3n}+\varepsilon}.$
Corollary 3.2.
Let $\varepsilon>0$ and let $q$ be a prime number. Then
$h_{2}(\mathbb{Q}[\sqrt[n]{q}])\ll_{n,\varepsilon}D_{\mathbb{Q}[\sqrt[n]{q}]}^{\frac{1}{2}-\frac{2}{3n}+\frac{1}{3n(n-1)}+\varepsilon}$
Proof.
One has $\|\alpha_{2,K}\|\ll q^{\frac{1}{n}}$ and $D_{\mathbb{Q}[\sqrt[n]{q}]}=q^{n-1}$. Then,
$\|\alpha_{2,K}\|\ll q^{\frac{1}{n}}=(q^{n-1})^{\frac{1}{n(n-1)}}=D_{\mathbb{Q}[\sqrt[n]{q}]}^{\frac{1}{n(n-1)}},$
which implies $\delta_{\mathbb{Q}[\sqrt[n]{q}],2}\leq\frac{1}{n(n-1)}$. Hence, an application of Theorem 2.3 complete the proof of the Corollary. ∎
4. Proof of Theorem 1.1
4.1. Some Lemmas
As mentioned in the introduction, the idea is to deploy the square sieve to detect how often the norm
$N_{K}(X_{1}+X_{2}\alpha_{2,K}+z_{3}\alpha_{3,K}+\cdots+z_{n}\alpha_{n,K})=:F_{\operatorname{\mathbf{z}}}(X_{1},X_{2}),$
is a square, for $(x_{1},x_{2})\in\left[-D_{K}^{\frac{1}{n}},D_{K}^{\frac{1}{n}}\right]\times\left[-D_{K}^{\frac{1}{n}-\delta_{2,K}},D_{K}^{\frac{1}{n}-\delta_{2,K}}\right]\cap\mathbb{Z}^{2}$, and where
$\operatorname{\mathbf{z}}\in\left[-D_{K}^{\frac{1}{n}-\delta_{3,K}},D_{K}^{\frac{1}{n}-\delta_{3,K}}\right]\times\cdots\times\left[-D_{K}^{\frac{1}{n}-\delta_{n,K}},D_{K}^{\frac{1}{n}-\delta_{n,K}}\right]\cap\mathbb{Z}^{n-2}$
is fixed. Before applying the sieve, we must exclude certain choices of $\operatorname{\mathbf{z}}$ for which our method does not apply. Namely:
  • $(i)$
    Those $\operatorname{\mathbf{z}}\in\mathbb{Z}^{n-2}$ such that the polynomial $N_{K}(X_{1}+X_{2}\alpha_{2,K}+z_{3}\alpha_{3,K}+\cdots+z_{n}\alpha_{n,K})$ is a square, i.e., the equation
    $Y^{2}=N_{K}(X_{1}+X_{2}\alpha_{2,K}+z_{3}\alpha_{3,K}+\cdots+z_{n}\alpha_{n,K})$
    has a solution for every choice of $X_{1},X_{2}$. We denote the set of these points as $Z_{\square}$.
  • $(ii)$
    Those $\operatorname{\mathbf{z}}\in\mathbb{Z}^{n-2}$ for which the the polynomial $F_{\operatorname{\mathbf{z}}}(X_{1},X_{2})$ does not satisfies the hypotheses of Proposition 2.4 over $\overline{\mathbb{Q}}$. We denote the set of these points as $Z_{F}$.
To simplify the notation, from now on we denote
$\mathcal{B}=\left[-D_{K}^{\frac{1}{n}},D_{K}^{\frac{1}{n}}\right]\times\left[-D_{K}^{\frac{1}{n}-\delta_{2,K}},D_{K}^{\frac{1}{n}-\delta_{2,K}}\right],$
and
$\mathcal{C}=\left[-D_{K}^{\frac{1}{n}-\delta_{3,K}},D_{K}^{\frac{1}{n}-\delta_{3,K}}\right]\times\cdots\times\left[-D_{K}^{\frac{1}{n}-\delta_{n,K}},D_{K}^{\frac{1}{n}-\delta_{n,K}}\right].$
4.1.1. The contribution of $Z_{\square}$
We start by recording the following Lemma, which is proved in [BSTTZ20, Bottom of page 66] and will be useful in bounding the contribution of the $\operatorname{\mathbf{z}}\in Z_{\square}$:
Lemma 4.1.
Let $K/\mathbb{Q}$ be a number field and assume that there exists $\mathbb{Q}\subset F\subset K$, such that $[K:F]=2$. Then $h_{2}(K)\ll_{\varepsilon}D_{K}^{\frac{1}{4}+\varepsilon}$.
Corollary 4.2.
Assume that $K/\mathbb{Q}$ does not contain a subfield of index 22. Then $Z_{\square}$ is empty if $n=5,6,7$, and
$\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\cap Z_{\square}\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1\ll_{n,\varepsilon}D_{K}^{\frac{1}{4}+\frac{1}{n}+\varepsilon},$
for $n\geq 8$.
Proof.
If the polynomial $F_{\operatorname{\mathbf{z}}}(X_{1},X_{2})$ is a square, then $F_{\operatorname{\mathbf{z}}}(X_{1},x_{2})=N_{K\mathbb{Q}}(X_{1}+x_{2}\alpha_{2,K}+z_{3}x_{3}+\cdots+z_{n}\alpha_{n,K})$ is a square for every value of $x_{2}\in\mathbb{Z}$. On the other hand, since we are assuming that $K$ does not contain any subfield of index 22, one can find $(x_{2},\operatorname{\mathbf{z}})$ such that $F_{\operatorname{\mathbf{z}}}(X_{1},x_{2})=N_{K\mathbb{Q}}(X_{1}+x_{2}\alpha_{2,K}+z_{3}x_{3}+\cdots+z_{n}\alpha_{n,K})$ is a square only if there exists $F\subset K$ such that $[K:F]>2$ even (see [BSTTZ20, proof of Lemma 4.1]). Hence, for $n=5,6,7$ the set $Z_{\square}$ is actually empty. Let us assume now $n\geq 8$. We start by writing
$\begin{split}\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\cap Z_{\square}\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1&=\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\cap\mathbb{Z}^{n-2}\\ F_{\operatorname{\mathbf{z}}}(X_{1},X_{2})=\square\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1\\ &\leq\sum_{\begin{subarray}{c}(x_{2},\operatorname{\mathbf{z}})\in\mathbb{Z}^{n-1}\\ (x_{2},\operatorname{\mathbf{z}})\in[-D_{K}^{\frac{1}{n}+\delta_{2,K}},D_{K}^{\frac{1}{n}+\delta_{2,K}}]\times\mathcal{C}\\ F_{\operatorname{\mathbf{z}}}(X_{1},x_{2})=\square\end{subarray}}\quad\sum_{\begin{subarray}{c}|x_{1}|\leq D_{K}^{\frac{1}{n}}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1.\end{split}$
An application of [BSTTZ20, Lemma 4.1] leads to
$\begin{split}\sum_{\begin{subarray}{c}(x_{2},\operatorname{\mathbf{z}})\in\mathbb{Z}^{n-1}\\ (x_{2},\operatorname{\mathbf{z}})\in[-D_{K}^{\frac{1}{n}+\delta_{2,K}},D_{K}^{\frac{1}{n}+\delta_{2,K}}]\times\mathcal{C}\\ F_{\operatorname{\mathbf{z}}}(X_{1},x_{2})=\square\end{subarray}}1\ll_{n,\varepsilon}D_{K}^{\frac{1}{4}+\varepsilon}\end{split}$
Hence
$\sum_{\begin{subarray}{c}(x_{2},\operatorname{\mathbf{z}})\in\mathbb{Z}^{n-1}\\ (x_{2},\operatorname{\mathbf{z}})\in[-D_{K}^{\frac{1}{n}+\delta_{2,K}},D_{K}^{\frac{1}{n}+\delta_{2,K}}]\times\mathcal{C}\\ F_{\operatorname{\mathbf{z}}}(X_{1},x_{2})=\square\end{subarray}}\quad\sum_{\begin{subarray}{c}|x_{1}|\leq D_{K}^{\frac{1}{n}}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1\ll_{n,\varepsilon}D_{K}^{\frac{1}{4}+\frac{1}{n}+\varepsilon},$
as claimed. ∎
4.1.2. The contribution of $Z_{F}$
We now bound the contribution of the vectors in $\mathcal{C}\cap Z_{F}$. Since we have
$\begin{split}F_{\operatorname{\mathbf{z}}}[X_{1},X_{2}]&=\quad\prod_{i=1}^{n}(X_{1}+\alpha_{2,K}X_{2}+z)\text{ over }\overline{\mathbb{Q}}\\ &=\prod_{\sigma\in\operatorname{Emb}(K)}(X_{1}+\sigma(\alpha_{2,K})X_{2}+\sigma(z))\text{ over }\overline{\mathbb{Q}},\end{split}$
the polynomial $F_{\operatorname{\mathbf{z}}}[X_{1},X_{2}]$ satisfies the hypothesis of Proposition 2.4 over $\overline{\mathbb{Q}}$, if
  1. $a)$
    the element $z$ does not belong in any intermediate field $\mathbb{Q}\subset F\subset K$ (this will also guarantees that $F_{\operatorname{\mathbf{z}}}$ is square-free);
  2. $b)$
    for every $\sigma,\tau\in\operatorname{Emb}(K)$, with $\sigma\neq\tau$, and $\tau(\alpha_{2,K})\neq\alpha$, one has
    $\sigma(z)-z\neq\frac{\sigma(\alpha_{2,K})-\alpha_{2,K}}{\tau(\alpha_{2,K})-\alpha_{2,K}}(\tau(z)-z).$
Notice that, since $\alpha_{2}\in K\setminus\mathbb{Q}$, the polynomial $F$ always satisfies condition $(i)$ of Proposition 2.4. We start with the following:
Lemma 4.3.
Let $\lambda\in K\setminus\{0,1\}$, and let $\sigma,\tau\in\operatorname{Emb}(K/\mathbb{Q})$ such that $\sigma\neq\tau$. Let $V_{\sigma,\tau,\lambda}$ be the $\mathbb{Q}$-vector space defined as
$V_{\sigma,\tau,\lambda}=\left\{x\in K:x=\frac{1}{1-\lambda}\sigma(x)-\frac{\lambda}{1-\lambda}\tau(x)\right\}.$
Then $\dim_{\mathbb{Q}}(V_{\sigma,\tau,\lambda})\leq\frac{2n}{3}$.
Proof.
Let $L$ be the Galois closure of $K$, observe that we can extend any embedding $\operatorname{Emb}(K/\mathbb{Q})$ to an element of $\operatorname{Gal}(L/\mathbb{Q})$: for every $\rho\in\operatorname{Emb}(K/\mathbb{Q})$, we will denote $\tilde{\rho}$ one of its extension to the whole $L$. Consider the linear map
$\varphi:K\rightarrow L,$
defined as $\varphi(x)=x-\frac{1}{1-\lambda}\sigma(x)+\frac{\lambda}{1-\lambda}\tau(x)$. Then $V_{\sigma,\tau,\lambda}=\text{Ker}(\varphi)$. Moreover, denote by $W_{\sigma,\tau,\lambda}=\text{Im}(\varphi)$. Let $d=\dim V_{\sigma,\tau,\lambda}$. Then $\dim_{\mathbb{Q}}W_{\sigma,\tau,\lambda}=n-d$, and let $\{h_{1},...,h_{n-d}\}$ be a basis of for $W_{\sigma,\tau,\lambda}$. Let $\rho_{1},...,\rho_{n-d+1}\in\operatorname{Emb}(K/\mathbb{Q})$ distinct morphisms and consider the system
[…]
Since this system has ndn-d equations in nd+1n-d+1 variables, one of the columns has to be a linear combination of the others. Hence, without loss of generality:
$\begin{pmatrix}\tilde{\rho}_{n-d+1}(h_{1})\\ \tilde{\rho}_{n-d+1}(h_{2})\\ \vdots\\ \tilde{\rho}_{n-d+1}(h_{n-d})\end{pmatrix}=\mu_{1}\begin{pmatrix}\tilde{\rho}_{1}(h_{1})\\ \tilde{\rho}_{1}(h_{2})\\ \vdots\\ \tilde{\rho}_{1}(h_{n-d})\end{pmatrix}+\cdots+\mu_{n-d}\begin{pmatrix}\tilde{\rho}_{n-d}(h_{1})\\ \tilde{\rho}_{n-d}(h_{2})\\ \vdots\\ \tilde{\rho}_{n-d}(h_{n-d}),\end{pmatrix}$
for some $\mu_{1},...,\mu_{n-d}$. Knowing that $h_{1},...,h_{n-d}$ is a basis for $W_{\sigma,\tau,\lambda}$ and that $\varphi$ vanishes on $V_{\sigma,\tau,\lambda}$, one has
$\tilde{\rho}_{n-d+1}\circ\varphi=\mu_{1}(\tilde{\rho}_{1}\circ\varphi)+\cdots\mu_{n-d}(\tilde{\rho}_{n-d}\circ\varphi).$
By contradiction, assume that $d>\frac{2n}{3}$. Then, $n-d<\frac{n}{3}$. We claim that for every $k<\frac{n}{3}$, we can find $A_{k}=\{\text{Id}=\rho_{1},\rho_{2},...,\rho_{k+1}\}\subset\operatorname{Emb}(K/\mathbb{Q})$, such that:
$\rho_{i}\neq\rho_{j},\quad\rho_{i}\neq\tilde{\rho}_{j}\circ\sigma,\quad\rho_{i}\neq\tilde{\rho}_{j}\circ\tau\quad\text{for any }j<i.$ (5)
We do this by induction, when $k=0$, we can take $\{\text{Id}\}$. Assume the claim is true for k1k-1, then the set
$B_{k-1}=\bigcup_{\rho\in A_{k-1}}\{\rho,\tilde{\rho}\circ\sigma,\tilde{\rho}\circ\tau\},$
has cardinality $B_{k-1}\leq 3k<n$ (we are assuming $k<\frac{n}{3}$). Since $\operatorname{Emb}(K/\mathbb{Q})=n$, we can find $\rho_{k+1}\in\operatorname{Emb}(K/\mathbb{Q})\setminus A_{k}$ which proves the claim. Since $n-d<\frac{n}{3}$, the claim implies that there exists $A_{n-d}=\{\text{Id}=\rho_{1},...,\rho_{n-d+1}\}\subset\operatorname{Emb}(K/\mathbb{Q})$ such that for any i=1,..,nd+1i=1,..,n-d+1
$\rho_{i}\neq\rho_{j},\quad\rho_{i}\neq\tilde{\rho}_{j}\circ\sigma,\quad\rho_{i}\neq\tilde{\rho}_{j}\circ\tau\quad\text{for any }j<i.$
Since we have nd+1n-d+1 endomorphisms, by the previous part (reordering the morphisms if necessary), we have that
$\tilde{\rho}_{n-d+1}\circ\varphi=\mu_{1}(\tilde{\rho}_{1}\circ\varphi)+\cdots\mu_{n-d}(\tilde{\rho}_{n-d}\circ\varphi).$
On the other hand, using the definition of $\varphi$, one would get that
$\begin{split}\rho_{n-d+1}&=-\frac{1}{1-\rho_{n-d+1}(\lambda)}(\tilde{\rho}_{n-d+1}\circ\sigma)+\frac{\rho_{n-d+1}(\lambda)}{1-\rho_{n-d+1}(\lambda)}(\tilde{\rho}_{i}\circ\tau)\\ &\quad+\sum_{i=1}^{n-d}\mu_{i}\left(\rho_{i}-\frac{1}{1-\rho_{i}(\lambda)}(\tilde{\rho}_{i}\circ\sigma)+\frac{\rho_{i}(\lambda)}{1-\rho_{i}(\lambda)}(\tilde{\rho}_{i}\circ\tau)\right),\end{split}$ (6)
[…]
We recall that, for every $\operatorname{\mathbf{z}}\in\mathbb{Z}^{n-2}$, we have $z=z_{3}\alpha_{3,K}+\cdots+z_{n}\alpha_{n,K}\in K$. Moreover, in what follow we denote
$E_{K,3}=\sum_{3\leq j\leq\frac{2n}{3}}\left(\frac{1}{n}-\delta_{j,K}\right).$
Corollary $4.$
Let $K/\mathbb{Q}$ be a number field of degree $n\geq 5$ which does not contain any subfield of index 22. Let $\sigma,\tau\in\operatorname{Emb}(K/\mathbb{Q})$ be two distinct element in $\operatorname{Emb}(K/\mathbb{Q})$ such that $\tau(\alpha_{2,K})\neq\alpha_{2,K}$. Then
$\left\{\operatorname{\mathbf{z}}\in\mathcal{C}\cap\mathbb{Z}^{n-2}:\sigma(z)-z=\frac{\sigma(\alpha_{2,K})-\alpha_{2,K}}{\tau(\alpha_{2,K})-\alpha_{2,K}}\cdot(\tau(z)-z)\right\}\leq D_{K}^{E_{K,3}}$
Proof.
Consider the vector space
$V_{\sigma,\tau}=\left\{x\in K:\sigma(x)-x=\frac{\sigma(\alpha_{2,K})-\alpha_{2,K}}{\tau(\alpha_{2,K})-\alpha_{2,K}}\cdot(\tau(x)-x)\right\}.$
We first handle the case when $\frac{\sigma(\alpha_{2,K})-\alpha_{2,K}}{\tau(\alpha_{2,K})-\alpha_{2,K}}=1$. In this case, $x$ satisfies $(\tau^{-1}\circ\sigma)(x)=x$, i.e. $x\in K^{\tau^{-1}\circ\sigma}$ which is a $\mathbb{Q}$-vectorspace of dimension at most $\frac{n}{3}$ (since we are assuming that $K$ does not contain any subfield of index 22). Since $1\in V_{\sigma,\tau}$, we have
$\dim\left(\left\{(z_{3},...,z_{n}):z_{3}\alpha_{3,K}+\cdots+z_{n}\alpha_{n,K}\in V_{\sigma,\tau}\right\}\right)\leq\frac{n}{3}-1.$
Thus,
$\{\operatorname{\mathbf{z}}\in\mathbb{C}\cap\mathbb{Z}^{n-2}:z\in V_{\sigma,\tau}\}\leq D_{K}^{\sum_{3\leq j\leq\frac{n}{3}+1}\left(\frac{1}{n}-\delta_{j,K}\right)}\leq D_{K}^{E_{K,3}},$
since when $n\geq 5$, $\frac{n}{3}+1\leq\frac{2n}{3}$. We now assume $\frac{\sigma(\alpha_{2,K})-\alpha_{2,K}}{\tau(\alpha_{2,K})-\alpha_{2,K}}\neq 1$. In this case, we can apply Lemma 4.3 with $\lambda=\frac{\sigma(\alpha_{2,K})-\alpha_{2,K}}{\tau(\alpha_{2,K})-\alpha_{2,K}}$, getting $\dim_{\mathbb{Q}}V\leq\frac{2n}{3}$. Moreover, one has that $1,\alpha_{2,K}\in V_{\sigma,\tau}$, thus
$\dim\left(\left\{(x_{3},...,x_{n}):x_{3}\alpha_{3,K}+\cdots+x_{n}\alpha_{n,K}\in V\right\}\right)=\frac{2n}{3}-2.$
Thus, it follows that
$(V_{\sigma,\tau}\cap\mathcal{C}\cap\mathbb{Z}^{n-2})\leq D_{K}^{E_{K,3}},$
as claimed. ∎
Lemma 4.5.
Some notation as Corollary 4.13. Then
$\{\operatorname{\mathbf{z}}\in\mathcal{C}\cap\mathbb{Z}^{n-2}:\sigma(z)=z,\text{ for some }\sigma\in\operatorname{Emb}(K/\mathbb{Q})\setminus\{\operatorname{Id}\}\}\ll D_{K}^{E_{K,3}}$
Proof.
One has that
$\{x\in K:\sigma(x)=x,\text{ for some }\sigma\in\operatorname{Emb}(K/\mathbb{Q})\setminus\{\operatorname{Id}\}\}=\bigcup_{\begin{subarray}{c}\sigma\in\operatorname{Emb}(K/\mathbb{Q})\\ \sigma\neq\text{Id}\end{subarray}}K^{\sigma},$
where $K^{\sigma}$ is the field of all the element in $K$ which are fixed by $\sigma$. Since $K:K^{\sigma}]\geq 3$ (we are assuming that $K$ does not contain any intermediate field of index 22), then $K^{\sigma}$ is a $\mathbb{Q}$-vector field of dimension $\leq\frac{n}{3}$, for every $\sigma\in\operatorname{Emb}(K/\mathbb{Q})\setminus\{\operatorname{Id}\}$. Moreover, since $1\in K^{\sigma}$
$\dim\left(\left\{(z_{3},...,z_{n}):z_{3}\alpha_{3,K}+\cdots+z_{n}\alpha_{n,K}\in K^{\sigma}\right\}\right)\leq\frac{n}{3}-1.$
Hence,
$\begin{split}\#\{\operatorname{\mathbf{z}}\in\mathcal{C}\cap\mathbb{Z}^{n-2}:\sigma(z)=z,\text{ for some }\sigma\in\operatorname{Emb}(K/\mathbb{Q})\setminus\{\operatorname{Id}\}\}&\ll D_{K}^{\sum_{3\leq j\leq\frac{n}{3}+1}\left(\frac{1}{n}-\delta_{j,K}\right)}\\ &\ll D_{K}^{\sum_{3\leq j\leq\frac{2n}{3}}\left(\frac{1}{n}-\delta_{j,K}\right)},\end{split}$
since when $n\geq 5$, $\frac{n}{3}+1\leq\frac{2n}{3}$
We are finally ready to prove
Corollary 4.6.
Assume that $K/\mathbb{Q}$ does not contain a subfield of index 22, and $\delta_{2,K}\geq\frac{1}{4n}$. Then
$\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\\ \operatorname{\mathbf{z}}\in Z_{F}\setminus Z_{\square}\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1\ll_{n,\varepsilon}D_{K}^{\frac{1}{2}-\frac{1}{2n}-\frac{1}{7n}}$
Proof.
We have:
$Z_{F}=\left(\bigcup_{\begin{subarray}{c}\sigma,\tau\in\operatorname{Emb}(K/\mathbb{Q})\\ \sigma\neq\tau\\ \tau(\alpha_{2,K})\neq\alpha_{2,K}\end{subarray}}V_{\sigma,\tau}\right)\cup\left(\bigcup_{\begin{subarray}{c}\sigma\in\operatorname{Emb}(K/\mathbb{Q})\\ \sigma\neq\operatorname{Id}\end{subarray}}K^{\sigma}\right)\cap\mathbb{Z}^{n-2},$
where
$V_{\sigma,\tau}=\left\{\operatorname{\mathbf{z}}\in\mathcal{C}:\sigma(z)-z=\frac{\sigma(\alpha_{2,K})-\alpha_{2,K}}{\tau(\alpha_{2,K})-\alpha_{2,K}}\cdot(\tau(z)-z)\right\}.$
Moreover, we recall that
$E_{K,3}=\sum_{3\leq j\leq\frac{2n}{3}}\left(\frac{1}{n}-\delta_{j,K}\right).$
$\begin{split}\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\\ \operatorname{\mathbf{z}}\in Z_{F}\setminus Z_{\square}\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1&=\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\cap Z_{F}\\ F_{\operatorname{\mathbf{z}}}(X_{1},X_{2})\neq\square\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1\\ &=\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\cap Z_{F}\\ F_{\operatorname{\mathbf{z}}}(X_{1},X_{2})\neq\square\end{subarray}}\quad\sum_{|x_{2}|\leq D_{K}^{\frac{1}{n}-\delta_{2,K}}}\sum_{\begin{subarray}{c}|x_{1}|\leq D^{\frac{1}{n}}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1\\ &\ll_{n,\varepsilon}D_{K}^{\frac{1}{2n}}\cdot D_{K}^{\frac{1}{n}-\delta_{2,K}}\cdot\#(\mathcal{C}\cap Z_{F})\end{split}$
where in the last step, we used the Bombieri-Pila bound ([BP89]). Using Corollary 4.4 and Lemma 4.5, one gets
$(\mathcal{C}\cap Z_{F})\ll D_{K}^{E_{K,3}}.$
[…]
$D_{K}^{\frac{1}{2n}+\left(\frac{1}{n}-\delta_{2,K}\right)+E_{K,3}}=D_{K}^{\frac{1}{2n}+\sum_{2\leq j\leq\frac{2n}{3}}\left(\frac{1}{n}-\delta_{j,K}\right)}\ll D_{K}^{\frac{1}{2}-\frac{1}{2n}-\sum_{j>\frac{2n}{3}}\left(\frac{1}{n}-\delta_{j,K}\right)}.$
If
$\sum_{j>\frac{2n}{3}-2}\left(\frac{1}{n}-\delta_{j,K}\right)<\frac{1}{7n},$
then
$\sum_{j>\frac{2n}{3}-2}\delta_{j,K}<\frac{1}{3}+\frac{1}{n}-\frac{1}{7n}.$
[…]
$\sum_{j=2}^{n}\delta_{j,K}=\frac{1}{2}+o(1);$
This would imply that
$\left\lfloor\frac{2n}{3}-2\right\rfloor\delta_{2,K}\leq\sum_{j=2}^{\lfloor\frac{2n}{3}-2\rfloor}\delta_{j,K}=\frac{1}{2}-\sum_{j>\frac{2n}{3}-2}\delta_{j,K}<\frac{1}{2}-\frac{1}{3}-\frac{1}{n}+\frac{1}{7n}\leq\frac{1}{6}-\frac{6}{7n}$
Thus, one deduces that
$\delta_{2,K}<\left(\frac{1}{6}-\frac{6}{7n}\right)\cdot\left(\frac{2n}{3}-3\right)^{-1}\leq\frac{1}{4n},\quad\text{ when $n\geq 5$}$
which would be in contradiction with the fact that we are assuming $\delta_{2,K}\geq\frac{1}{4n}$. Hence, we have
$\sum_{j>\frac{2n}{3}-2}\left(\frac{1}{n}-\delta_{j,K}\right)\geq\frac{1}{7n},$
which leads to
$\begin{split}\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\cap Z_{F}\\ F_{\operatorname{\mathbf{z}}}(X_{1},X_{2})\neq\square\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1\ll_{n,\varepsilon}D_{K}^{\frac{1}{2}-\frac{1}{2n}-\frac{1}{7n}},\end{split}$
which conclude the proof of the Lemma. ∎
4.2. Preparation for the sieve: smooth weights.
As previously mentioned, our strategy is to fix $\operatorname{\mathbf{z}}=(z_{3},\dots,z_{n})$ and study how often
$F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=N_{K}\left(X_{1}+X_{2}\alpha_{2,K}+z_{3}\alpha_{3,K}+\cdots+z_{n}\alpha_{n,K}\right)$
is a square, as $(x_{1},x_{2})$ ranges over the box
$\mathcal{B}_{\operatorname{\mathbf{z}}}:=\left[-D_{K}^{\frac{1}{n}},D_{K}^{\frac{1}{n}}\right]\times\left[-D_{K}^{\frac{1}{n}-\delta_{2,K}},D_{K}^{\frac{1}{n}-\delta_{2,K}}\right].$
Having this in mind, define
$a_{\operatorname{\mathbf{z}}}(n)=\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}_{\operatorname{\mathbf{z}}}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=n\end{subarray}}1.$
Then,
$\left\{(x_{1},x_{2})\in\mathcal{B}_{\operatorname{\mathbf{z}}}:F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\right\}\leq\sum_{k\in\mathbb{N}}a_{\operatorname{\mathbf{z}}}(k^{2}).$ (7)
Our goal is to estimate the right-hand side of (7) using the square sieve (including its version for composite moduli). As we will see in the upcoming sections, after applying the sieve lemma, we will need to apply the Poisson summation formula. For that reason, rather than working directly with the sequence $(a_{\operatorname{\mathbf{z}}}(n))_{n\in\mathbb{N}}$, we will instead introduce suitable smooth weights. We begin by choosing two smooth functions $g_{1},g_{2}:\mathbb{R}\to\mathbb{R}_{\geq 0}$ satisfying the following properties:
  • $i)$
    $g_{1}$ is supported on $\left[-D_{K}^{\frac{1}{n}},D_{K}^{\frac{1}{n}}\right]$ and $g_{2}$ is supported on $\left[-D_{K}^{\frac{1}{n}-\delta_{2,K}},D_{K}^{\frac{1}{n}-\delta_{2,K}}\right]$
  • ii)ii)
    for any $k\geq 0$, we have that
    $\left|\frac{\partial^{k}g_{1}}{\partial x_{1}^{k}}\right|\ll D_{K}^{-\frac{k}{n}},\quad\left|\frac{\partial^{k}g_{2}}{\partial x_{2}^{k}}\right|\ll D_{K}^{-k\left(\frac{1}{n}-\delta_{2,K}\right)}.$ (8)
From the condition above, one deduces the following bounds on the Fourier transforms of $g_{1}$, and $g_{2}$
$\hat{g}_{1}(u)|\ll D_{K}^{\frac{1}{n}}\left|1+uD_{K}^{\frac{1}{n}}\right|^{-2},\quad|\hat{g}_{2}(u)|\ll D_{K}^{\frac{1}{n}-\delta_{2,K}}\left|1+uD_{K}^{\frac{1}{n}-\delta_{2,K}}\right|^{-2}.$
Then, for any $n\geq 0$ we define
$\omega_{\operatorname{\mathbf{z}}}(n)=\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=n\end{subarray}}g_{1}(x_{1})g_{2}(x_{2}).$ (9)
We remark that when $n=0$,
$\omega_{\operatorname{\mathbf{z}}}(n)=\begin{cases}g_{1}(0)g_{2}(0)&\text{if $\operatorname{\mathbf{z}}=\boldsymbol{0}$}\\ 0&\text{otherwise,}\end{cases}$
since $F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=N_{K}(x_{1}+x_{2}\alpha_{2,K}+z_{3}\alpha_{3,K}+\cdots+z_{n}\alpha_{n,K})$ and $x\in K$ is such that $N_{K}(x)=0$ if and only if $x=0$.
4.3. The square sieve: proof of Theorem 2.3.
We record the following Sieve Lemma due to Heath-Brown [HB84]:
Lemma 4.7.
Let $\mathcal{P}$ be a finite set of $P$ prime numbers. Let $(\omega(n))_{n\in\mathbb{N}}$ be a sequence of non negative number such that $\omega(n)=0$ if either $n=0$, or $n>e^{P}$. Then
$\sum_{k}\omega(k^{2})\ll\frac{1}{P}\sum_{n}\omega(n)+\frac{1}{P^{2}}\sum_{\begin{subarray}{c}p,q\in\mathcal{P}\\ p\neq q\end{subarray}}\left|\sum_{n}\omega(n)\left(\frac{n}{pq}\right)\right|.$
We will use the sieve inequality in Lemma 4.7 to show:
Lemma 4.8.
[…]
$\sum_{k}\omega_{\operatorname{\mathbf{z}}}(k^{2})\ll_{n,\varepsilon}D_{K}^{\frac{4}{3n}-\frac{2\delta_{2,K}}{3}+\varepsilon}.$
To prove Lemma 4.8, we start by applying Lemma 4.7 to the sequence $(\omega_{\operatorname{\mathbf{z}}}(n))_{n}$:
$\sum_{k}\omega_{\operatorname{\mathbf{z}}}(k^{2})\ll\frac{1}{P}\sum_{n}\omega_{\operatorname{\mathbf{z}}}(n)+\frac{1}{P^{2}}\sum_{\begin{subarray}{c}p,q\in\mathcal{P}\\ p\neq q\end{subarray}}\left|\sum_{n}\omega_{\operatorname{\mathbf{z}}}(n)\left(\frac{n}{pq}\right)\right|,$ (10)
provided that $P>(\log D_{K})^{2}$. By the definition of the sequence $\omega_{\operatorname{\mathbf{z}}}(n)$, it follows that
$\frac{1}{P}\sum_{n}\omega_{\operatorname{\mathbf{z}}}(n)\ll\frac{D_{K}^{\frac{2}{n}-\delta_{2,K}}}{P}.$ (11)
[…]
$\begin{split}S(\operatorname{\mathbf{z}},p,q)&=\sum_{n\in\mathbb{N}}\omega_{\operatorname{\mathbf{z}}}(n)\left(\frac{n}{pq}\right)\\ &=\sum_{(x_{1},x_{2})\in\mathbb{Z}^{2}}g_{1}(x_{1})g_{2}(x_{2})\left(\frac{F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})}{pq}\right)\\ &=\sum_{(a_{1},a_{2})\text{ }(pq)}\left(\frac{F_{\operatorname{\mathbf{z}}}(a_{1},a_{2})}{pq}\right)\sum_{\begin{subarray}{c}x_{1}\in\mathbb{Z}\\ x_{1}\equiv a_{1}\text{ }(pq)\end{subarray}}g_{1}(x_{1})\sum_{\begin{subarray}{c}x_{2}\in\mathbb{Z}\\ x_{2}\equiv a_{2}\text{ }(pq)\end{subarray}}g_{2}(x_{2})\\ &=\frac{1}{(pq)^{2}}\sum_{(u_{1},u_{2})\in\mathbb{Z}^{2}}\hat{g}_{1}\left(\frac{u_{1}}{pq}\right)\hat{g}_{2}\left(\frac{u_{2}}{pq}\right)T(\operatorname{\mathbf{u}},\operatorname{\mathbf{z}};pq),\end{split}$ (12)
where for every $r\geq 2$ square-free and $\operatorname{\mathbf{v}}\in\mathbb{Z}^{2}$,
$T(\operatorname{\mathbf{v}},\operatorname{\mathbf{z}};r)=\sum_{(a_{1},a_{2})\text{ }(r)}\left(\frac{F_{\operatorname{\mathbf{z}}}(a_{1},a_{2})}{r}\right)e_{r}(a_{1}v_{1}+a_{2}v_{2}).$
An application of the Chinese Remainder Theorem shows that
$S(\operatorname{\mathbf{z}},p,q)=\frac{1}{(pq)^{2}}\sum_{(u_{1},u_{2})\in\mathbb{Z}^{2}}\hat{g}_{1}\left(\frac{u_{1}}{pq}\right)\hat{g}_{2}\left(\frac{u_{2}}{pq}\right)T(\overline{q}\operatorname{\mathbf{u}},\operatorname{\mathbf{z}};p)T(\overline{p}\operatorname{\mathbf{u}},\operatorname{\mathbf{z}};q).$
4.3.1. The exponential sum $T(\operatorname{\mathbf{v}},\operatorname{\mathbf{z}};r)$.
In the next session (see Lemma 5.6), we shall prove the following
Lemma 4.9.
Let $\operatorname{\mathbf{z}}=(z_{3},...,z_{n})\in\mathbb{Z}^{n-2}$, such that $F_{\operatorname{\mathbf{z}}}(X_{1},X_{2})$ is not a square. Then for every prime number $p$ such that
  • $i)$
    the polynomial $F_{\operatorname{\mathbf{z}}}(X_{1},X_{2})\mod p$ is not a square;
  • ii)ii)
    $p\nmid N_{K}(\sigma(\alpha_{2})-\alpha_{2})$, for every $\sigma\in\operatorname{Emb}(K/\mathbb{Q})$ such that $\sigma(\alpha_{2})\neq\alpha_{2}$.
$|T(\operatorname{\mathbf{v}},\operatorname{\mathbf{z}};p)|\ll_{n}p(p,v_{1},v_{2})^{\frac{1}{2}}.$
As a consequence of Lemma 4.9, we obtain
Corollary 4.10.
Let $Q>0$ and let $p,q$ be primes such that $Q<p,q<2Q$ and assume
  • $i)$
    the polynomials $F_{\operatorname{\mathbf{z}}}(X_{1},X_{2})\mod p$, $F_{\operatorname{\mathbf{z}}}(X_{1},X_{2})\mod q$ are not squares in $\mathbb{F}_{p}$ and $\mathbb{F}_{q}$ respectively;
  • ii)ii)
    $p,q\nmid N_{K}(\sigma(\alpha_{2})-\alpha_{2})$, for every $\sigma\in\operatorname{Emb}(K/\mathbb{Q})$ such that $\sigma(\alpha_{2})\neq\alpha_{2}$.
Then
$S(\operatorname{\mathbf{z}},pq)\ll_{n}Q^{2}+\frac{D^{\frac{2}{n}-\delta_{2,K}}}{Q},$
provided that $D_{K}^{\frac{1}{2n}}\leq Q\leq D_{K}^{\frac{1}{n}-\delta_{2,K}}$.
Proof.
Since the bound for $T(\operatorname{\mathbf{u}},\operatorname{\mathbf{z}},pq)$ depends on the G.C.D. between $u_{1},u_{2}$ and pqpq we start by rewriting $S(\operatorname{\mathbf{z}},pq)$ as
$S(\operatorname{\mathbf{z}},pq)=\frac{1}{(pq)^{2}}\sum_{\begin{subarray}{c}\operatorname{\mathbf{u}}\in\mathbb{Z}^{2}\\ (\operatorname{\mathbf{u}},pq)=1\end{subarray}}\hat{g}_{1}\left(\frac{u_{1}}{pq}\right)\hat{g}_{2}\left(\frac{u_{2}}{pq}\right)T(\operatorname{\mathbf{u}},\operatorname{\mathbf{z}};pq)+\frac{1}{(pq)^{2}}\sum_{\begin{subarray}{c}\operatorname{\mathbf{u}}\in\mathbb{Z}^{2}\\ (\operatorname{\mathbf{u}},pq)>1\end{subarray}}\hat{g}_{1}\left(\frac{u_{1}}{pq}\right)\hat{g}_{2}\left(\frac{u_{2}}{pq}\right)T(\operatorname{\mathbf{u}},\operatorname{\mathbf{z}};pq)$
[…]
$\left|\hat{g}_{1}\left(\frac{u_{1}}{pq}\right)\hat{g}_{2}\left(\frac{u_{2}}{pq}\right)\right|\ll D_{K}^{\frac{2}{n}-\delta_{2,K}}\left(1+\frac{D_{K}^{\frac{1}{n}}u_{1}}{pq}\right)^{-2}\cdot\left(1+\frac{D_{K}^{\frac{1}{n}-\delta_{2,K}}u_{2}}{pq}\right)^{-2}.$
On the other hand, by applying Lemma 4.9, one gets
$\begin{split}\frac{1}{(pq)^{2}}\sum_{\begin{subarray}{c}\operatorname{\mathbf{u}}\in\mathbb{Z}^{2}\\ (\operatorname{\mathbf{u}},pq)=1\end{subarray}}|T(\operatorname{\mathbf{u}},\operatorname{\mathbf{z}};pq)|&\cdot\left|\hat{g}_{1}\left(\frac{u_{1}}{pq}\right)\hat{g}_{2}\left(\frac{u_{2}}{pq}\right)\right|\ll_{n}\\ &\ll_{n}pq\cdot\frac{D_{K}^{\frac{2}{n}-\delta_{2,K}}}{(pq)^{2}}\sum_{\begin{subarray}{c}\operatorname{\mathbf{u}}\in\mathbb{Z}^{2}\end{subarray}}\left(1+\frac{D_{K}^{\frac{1}{n}}u_{1}}{pq}\right)^{-2}\cdot\left(1+\frac{D_{K}^{\frac{1}{n}-\delta_{2,K}}u_{2}}{pq}\right)^{-2}\\ &\ll Q^{2},\end{split}$
provided that $D_{K}^{\frac{1}{n}}\leq pq\leq 4Q^{2}$. We now deal with the contributions of those $\operatorname{\mathbf{u}}$ for which $(\operatorname{\mathbf{u}},pq)>1$:
$\begin{split}\frac{1}{(pq)^{2}}\sum_{\begin{subarray}{c}\operatorname{\mathbf{u}}\in\mathbb{Z}^{2}\\ (\operatorname{\mathbf{u}},pq)>1\end{subarray}}\hat{g}_{1}\left(\frac{u_{1}}{pq}\right)\hat{g}_{2}\left(\frac{u_{2}}{pq}\right)T(\operatorname{\mathbf{u}},\operatorname{\mathbf{z}};pq)&\ll_{n}(pq)^{\frac{3}{2}}\cdot\frac{1}{(pq)^{2}}\sum_{\begin{subarray}{c}\operatorname{\mathbf{u}}\in\mathbb{Z}^{2}\\ p|\operatorname{\mathbf{u}}\end{subarray}}\hat{g}_{1}\left(\frac{u_{1}}{pq}\right)\hat{g}_{2}\left(\frac{u_{2}}{pq}\right)\\ &\quad+(pq)^{\frac{3}{2}}\cdot\frac{1}{(pq)^{2}}\sum_{\begin{subarray}{c}\operatorname{\mathbf{u}}\in\mathbb{Z}^{2}\\ q|\operatorname{\mathbf{u}}\end{subarray}}\hat{g}_{1}\left(\frac{u_{1}}{pq}\right)\hat{g}_{2}\left(\frac{u_{2}}{pq}\right).\end{split}$
Provided that $q\leq D^{\frac{1}{n}-\delta_{2,K}}$, one has
$\begin{split}\sum_{\begin{subarray}{c}\operatorname{\mathbf{u}}\in\mathbb{Z}^{2}\\ q|\operatorname{\mathbf{u}}\end{subarray}}\hat{g}_{1}\left(\frac{u_{1}}{pq}\right)\hat{g}_{2}\left(\frac{u_{2}}{pq}\right)&\ll D_{K}^{\frac{2}{n}-\delta_{2,K}}\sum_{\begin{subarray}{c}\operatorname{\mathbf{u}}\in\mathbb{Z}^{2}\\ q|\operatorname{\mathbf{u}}\end{subarray}}\left(1+\frac{D_{K}^{\frac{1}{n}}u_{1}}{pq}\right)^{-2}\cdot\left(1+\frac{D_{K}^{\frac{1}{n}-\delta_{2,K}}u_{2}}{pq}\right)^{-2}\\ &\ll D_{K}^{\frac{2}{n}-\delta_{2,K}},\end{split}$
which is enough to complete the proof of the Lemma. ∎
We are finally ready to complete the proof of Lemma 4.8: let $Q$ be such that $D_{K}^{\frac{1}{2n}}\leq Q\leq D_{K}^{\frac{1}{n}-\delta_{2,K}}$ and take
[…]
$\sum_{n\in\mathbb{Z}}\omega_{\operatorname{\mathbf{z}}}(k^{2})\ll_{n}\frac{D_{K}^{\frac{2}{n}-\delta_{2,K}}\log Q}{Q}+Q^{2}.$
The optimal choice here is when $Q=D_{K}^{\frac{2}{3n}-\frac{\delta_{2,K}}{3}}$. So we get
$\sum_{n\in\mathbb{Z}}\omega_{\operatorname{\mathbf{z}}}(k^{2})\ll_{n,\varepsilon}D_{K}^{\frac{4}{3n}-\frac{2\delta_{2,K}}{3}+\varepsilon},$
provided that $D_{K}^{\frac{1}{n}}\leq D_{K}^{\frac{4}{3n}-\frac{2\delta_{2,K}}{3}}$ and $D_{K}^{\frac{2}{3n}-\frac{\delta_{2,K}}{3}}\leq D^{\frac{1}{n}-\delta_{2,K}}$, which is satisfied as soon as $\delta_{2,K}\leq\frac{1}{2n}$.
4.3.2. Proof of Theorem 2.2
We can now prove Theorem $2.$ Thanks to Lemma 4.1, we may assume that $K$ does not contain a subfield of index 22. We recall that
$\mathcal{C}=\left[-D_{K}^{\frac{1}{n}-\delta_{3,K}},D_{K}^{\frac{1}{n}-\delta_{3,K}}\right]\times\cdots\times\left[-D_{K}^{\frac{1}{n}-\delta_{n,K}},D_{K}^{\frac{1}{n}-\delta_{n,K}}\right],$
and that
$Z_{\square}=\{\operatorname{\mathbf{z}}\in\mathbb{Z}^{n-2}:F_{\operatorname{\mathbf{z}}}(X_{1},X_{2})\text{ is a square}\}.$
Then
$\begin{split}\#h_{2}(K)&\ll\#\left\{(y,z_{1},...,z_{n})\in\mathbb{Z}^{n+1}:\begin{matrix}|z_{i}|\ll D_{K}^{\frac{1}{n}-\delta_{i,K}},\quad\text{for }i=1,...,n\\ y^{2}=N_{K}(z_{1}+z_{2}\alpha_{2,K}+\cdots+z_{n}\alpha_{n,K})\end{matrix}\right\}\\ &=\sum_{\operatorname{\mathbf{z}}\in\mathcal{C}\cap\mathbb{Z}^{n-2}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1\\ &=\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\cap Z_{\square}\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1+\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\cap\mathbb{Z}^{n-2}\\ \operatorname{\mathbf{z}}\not\in Z_{\square}\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1.\end{split}$ (13)
[…]
$\begin{split}\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\cap\mathbb{Z}^{n-2}\\ F_{\operatorname{\mathbf{z}}}(X_{1},X_{2})\neq\square\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1&\ll_{n,\varepsilon}\#(\mathcal{C}\cap\mathbb{Z})\cdot D_{K}^{\frac{4}{3n}-\frac{2\delta_{2,K}}{3}+\varepsilon}\\ &=D_{K}^{\sum_{j=3}^{n}\left(\frac{1}{n}-\delta_{j,K}\right)+\frac{4}{3n}-\frac{2\delta_{2,K}}{3}+\varepsilon}\\ &=D_{K}^{\frac{1}{2}-\frac{2}{3n}+\frac{\delta_{2,K}}{3}+\varepsilon}.\end{split}$
On the other, thanks to Lemma 4.2 we know that
$\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\cap Z_{\square}\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1\ll_{n,\varepsilon}D_{K}^{\frac{1}{4}+\frac{1}{n}+\varepsilon},$
when $n\geq 8$ and zero otherwise. Since $\frac{1}{4}+\frac{1}{n}\leq\frac{1}{2}-\frac{2}{3n}$, for $n>6$, we complete the proof of Theorem $2.$
$4.$ The square sieve for composite moduli: bounding $h_{2}(K)$ when $\delta_{2,K}$ is large.
In this section, we are going to prove Theorem 2.3. As mentioned in the introduction, we do this by deploying the square sieve for composite moduli. We record here the sieve Lemma as stated in [HP12, Lemma 22]
Lemma 4.11.
Let $A=|\mathcal{A}|$, $U=|\mathcal{U}|$ , and $V=|\mathcal{V}|$, so that $A=UV$ . Furthermore, assume that $V^{3}\ll A$. Let $\omega$ be a non-negative weight such that $\omega(n)=0$ for $n>\exp\min(U,V)$, then
$\begin{split}\sum_{n\neq 0}\omega(n)&\ll\frac{1}{A}\sum_{n\neq 0}\omega(n)+\frac{1}{A^{2}}\sum_{v,v^{\prime}\in\mathcal{V}}\sum_{\begin{subarray}{c}u,u^{\prime}\mathcal{U}\\ u\neq u^{\prime}\end{subarray}}\left|\sum_{n\in\mathbb{Z}}\left(\frac{n}{uu^{\prime}}\right)\left(\frac{n}{uu^{\prime}}\right)\omega(n)\right|\\ &\quad+\frac{U}{A^{2}}\sum_{\begin{subarray}{c}v,v^{\prime}\mathcal{V}\\ v\neq v^{\prime}\end{subarray}}\left|\sum_{n\in\mathbb{Z}}\left(\frac{n}{vv^{\prime}}\right)\omega(n)\right|.\end{split}$
In the remainder of this section we will use Lemma 4.11, to prove the following
Lemma 4.12.
Let $\operatorname{\mathbf{z}}\in\mathbb{Z}^{n-2}\setminus\{\boldsymbol{0}\}$ satisfying the following condition:
  • $i)$
    The polynomial $F_{\operatorname{\mathbf{z}}}(X_{1},X_{2})$ is not a square in $\mathbb{Z}[X_{1},X_{2}]$.
  • ii)ii)
    For every $\sigma,\tau\in\operatorname{Emb}(K/\mathbb{Q})$, with $\sigma\neq\tau$ and $\tau(\alpha_{2,K})\neq\alpha_{2,K}$,
    $\sigma(z)-z\neq\frac{\sigma(\alpha_{2,K})-\alpha_{2,K}}{\tau(\alpha_{2,K})-\alpha_{2,K}}(\tau(z)-z),$
  • $iii)$
    For every $\sigma\in\operatorname{Emb}(K/\mathbb{Q})\setminus\{\operatorname{Id}\}$, $z\not\in K^{\sigma}$.
[…]
$\sum_{k}\omega_{\operatorname{\mathbf{z}}}(k^{2})\ll_{n,\varepsilon}D_{K}^{\frac{45}{28n}-\frac{11}{14}\delta_{2,K}+\varepsilon}$
and ultimately, use Lemma 4.12 to prove Theorem 2.3.
To prove Lemma 4.12, we start by applying Lemma 4.11, getting
$\begin{split}\sum_{n\in\mathbb{Z}}\omega_{\operatorname{\mathbf{z}}}(n^{2})&=\frac{1}{A}\sum_{n}\omega_{\operatorname{\mathbf{z}}}(n)+\frac{1}{A^{2}}\sum_{v,v^{\prime}\in\mathcal{V}}\sum_{\begin{subarray}{c}u,u^{\prime}\mathcal{U}\\ u\neq u^{\prime}\end{subarray}}\left|\sum_{n\in\mathbb{Z}}\left(\frac{n}{uu^{\prime}}\right)\left(\frac{n}{uu^{\prime}}\right)\omega_{\operatorname{\mathbf{z}}}(n)\right|\\ &\quad+\frac{U}{A^{2}}\sum_{\begin{subarray}{c}v,v^{\prime}\mathcal{V}\\ v\neq v^{\prime}\end{subarray}}\left|\sum_{n\in\mathbb{Z}}\left(\frac{n}{vv^{\prime}}\right)\omega_{\operatorname{\mathbf{z}}}(n)\right|.\end{split}$ (14)
[…]
$\frac{1}{A}\sum_{n}\omega_{\operatorname{\mathbf{z}}}(n)\ll\frac{D_{K}^{\frac{2}{n}-\delta_{2,K}}}{A}.$ (15)
Hence, we need to focus on the two other terms. We will refer to the term
$M=\frac{1}{A^{2}}\sum_{v,v^{\prime}\in\mathcal{V}}\sum_{\begin{subarray}{c}u,u^{\prime}\mathcal{U}\\ u\neq u^{\prime}\end{subarray}}\left|\sum_{n\in\mathbb{Z}}\left(\frac{n}{uu^{\prime}}\right)\left(\frac{n}{uu^{\prime}}\right)\omega_{\operatorname{\mathbf{z}}}(n)\right|,$ (16)
as the main sieve and to the term
$N=\frac{U}{A^{2}}\sum_{\begin{subarray}{c}v,v^{\prime}\mathcal{V}\\ v\neq v^{\prime}\end{subarray}}\left|\sum_{n\in\mathbb{Z}}\left(\frac{n}{vv^{\prime}}\right)\omega_{\operatorname{\mathbf{z}}}(n)\right|,$
as the prime sieve.
Proof of Lemma 4.12: the main sieve
In what follows, we denote $q_{1}=uu^{\prime}$, $q_{2}=vv^{\prime}$. We start again from
$\begin{split}S(\operatorname{\mathbf{z}};q_{1}q_{2})&=\sum_{(x_{1},x_{2})\in\mathbb{Z}^{2}}g_{1}(x_{1})g_{2}(x_{2})\left(\frac{F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})}{q_{1}q_{2}}\right)\end{split}$
but this time, the first step will be to extend only the summation in $x_{1}$ over all the classes modulo $q_{1}q_{2}$:
$\begin{split}S(\operatorname{\mathbf{z}};q_{1}q_{2})&=\sum_{x_{2}\in\mathbb{Z}}g_{2}(x_{2})\sum_{y_{1}\text{ }(q_{1}q_{2})}\left(\frac{F_{\operatorname{\mathbf{z}}}(y_{1},x_{2})}{q_{1}q_{2}}\right)\sum_{\begin{subarray}{c}x_{2}\in\mathbb{Z}\\ x_{1}\equiv y_{1}\text{ }(q_{1}q_{2})\end{subarray}}g_{1}(x_{1})\\ &=\frac{1}{q_{1}q_{2}}\sum_{u_{1}}\hat{g}_{1}\left(\frac{u_{1}}{q_{1}q_{2}}\right)S(\operatorname{\mathbf{z}},u_{1};q_{1}q_{2}),\end{split}$
where
$S(\operatorname{\mathbf{z}},u_{1};q_{1}q_{2})=\sum_{x_{2}\in\mathbb{Z}}g_{2}(x_{2})Z(u_{1},x_{2};q_{1}q_{2}),$
and where for every $r$ square-free number,
$Z(u_{1},x_{2};r)=\sum_{y_{1}\text{ }(r)}\left(\frac{F_{\operatorname{\mathbf{z}}}(y_{1},x_{2})}{r}\right)e_{r}(u_{1}y_{1}).$
Again, applying the Chinese Remainder Theorem, one has that
$S(\operatorname{\mathbf{z}},u_{1};q_{1}q_{2})=\sum_{x_{2}\in\mathbb{Z}}g_{2}(x_{2})Z(u_{1}\overline{q_{2}},x_{2};q_{1})Z(u_{1}\overline{q_{1}},x_{2};q_{2})$
We now deploy the $q$-analogue of the Van der Corput method: we denote $H=\lfloor\frac{D_{K}^{\frac{1}{n}-\mu_{2}}}{q_{2}}\rfloor$. Then
$\begin{split}HS(\operatorname{\mathbf{z}},u_{1};q_{1}q_{2})&=\sum_{h\in[0,H]}\sum_{x_{2}\in\mathbb{Z}}g_{2}(x_{2}+q_{2}h)Z(u_{1}\overline{q_{2}},x_{2}+q_{2}h;q_{1})Z(u_{1}\overline{q_{1}},x_{2}+q_{2}h;q_{2})\\ &=\sum_{x_{2}\in I}Z(u_{1}\overline{q_{1}},x_{2};q_{2})\sum_{h\in[0,H]}g_{2}(x_{2}+q_{2}h)Z(u_{1}\overline{q_{2}},x_{2}+q_{2}h;q_{1}),\end{split}$
where $I=[D_{K}^{\frac{1}{n}-\delta_{2,K}}-q_{2}H,D_{K}^{\frac{1}{n}-\delta_{2,K}}-q_{2}]$. Applying Cauchy–Schwarz,
$H^{2}|S(\operatorname{\mathbf{z}},u_{1};q_{1}q_{2})|^{2}\ll\Sigma_{1}\Sigma_{2},$
where
$\begin{split}&\Sigma_{1}=\sum_{x_{2}\in I}|Z(u_{1}\overline{q_{1}},x_{2};q_{2})|^{2}\\ &\Sigma_{2}=\sum_{x_{2}\in I}\left|\sum_{h\in[0,H]}g_{2}(x_{2}+q_{2}h)Z(u_{1}\overline{q_{2}},x_{2}+q_{2}h;q_{1})\right|^{2}.\end{split}$
Expanding the square in $\Sigma_{2}$, we get
$\Sigma_{2}=\sum_{h_{1},h_{2}\in[0,H]}R(h_{1},h_{2}),$
where
$R(h_{1},h_{2})=\sum_{\begin{subarray}{c}x_{2}\in\mathbb{Z}\end{subarray}}g_{2}(x_{2}+q_{2}h_{1})g_{2}(x_{2}+q_{2}h_{2})Z(u_{1}\overline{q_{2}},x_{2}+q_{2}h_{1};q_{1})\overline{Z(u_{1}\overline{q_{2}},x_{2}+q_{2}h_{2};q_{1})}.$
Since $R(h_{1},h_{2})=R(h_{1}-h_{2},0)$, we get that
$\begin{split}\Sigma_{2}&=\sum_{h_{1},h_{2}\in[0,H]}R(h_{1}-h_{2},0)\\ &=\sum_{h\in[-H,H]}(H-|h|)R(h,0)\\ &\ll H\sum_{h\in[-H,H]}\left|\sum_{\begin{subarray}{c}x_{2}\in\mathbb{Z}\end{subarray}}g_{2}(x_{2})g_{2}(x_{2}+q_{2}h)Z(u_{1}\overline{q_{2}},x_{2};q_{1})\overline{Z(u_{1}\overline{q_{2}},x_{2}+q_{2}h;q_{1})}\right|,\end{split}$
Then
$\Sigma_{2}\leq H(\Sigma_{2A}+\Sigma_{2B}),$
where,
$\begin{split}&\Sigma_{2A}=\sum_{\begin{subarray}{c}x_{2}\in\mathbb{Z}\end{subarray}}(g(x_{2}))^{2}|Z(u_{1}\overline{q_{2}},x_{2};q_{1})|^{2}\\ &\Sigma_{2B}=\sum_{\begin{subarray}{c}h\in[-H,H]\\ h\neq 0\end{subarray}}\left|\sum_{\begin{subarray}{c}x_{2}\in\mathbb{Z}\end{subarray}}g_{2}(x_{2})g_{2}(x_{2}+q_{2}h)Z(u_{1}\overline{q_{2}},x_{2};q_{1})\overline{Z(u_{1}\overline{q_{2}},x_{2}+q_{2}h;q_{1})}\right|.\end{split}$
An application of Lemma 5.4 gives
$\Sigma_{1}\ll_{n}D_{K}^{\frac{1}{n}-\delta_{2,K}}q_{2},\quad\Sigma_{2A}\ll_{n}D_{K}^{\frac{1}{n}-\delta_{2,K}}q_{1},$
provided that $\mathcal{U},\mathcal{V}\subset\{p:F(X,Y)\mod p\text{ is not a square}\}.$ Thus, it remains to estimate $\Sigma_{2B}$. To do this, we first denote $g_{h}(x_{2})=g_{2}(x_{2})g_{2}(x_{2}+q_{2}h)$. Applying the Poisson summation formula, one gets
$\Sigma_{2B}\ll\frac{1}{q_{1}}\sum_{\begin{subarray}{c}h\in[-H,H]\end{subarray}}\left|\sum_{u_{2}\in\mathbb{Z}}\hat{g}_{h}\left(\frac{u_{2}}{q_{1}}\right)W(u_{1}\overline{q_{2}},u_{2},q_{2}h;q_{1})\right|$
where
$W(s,t,r;q_{1})=\sum_{x,y,w\text{ }(q_{1})}\left(\frac{F_{\operatorname{\mathbf{z}}}(x,y)}{q_{1}}\right)\left(\frac{F_{\operatorname{\mathbf{z}}}(w,y+r)}{q_{1}}\right)e_{p}(s(x-w)+ty)$
Applying Proposition 2.4, we have
Corollary 4.13.
Let $\operatorname{\mathbf{z}}=(z_{3},...,z_{n})$, and $q_{1}=uu^{\prime}$ with $u,u^{\prime}$ primes such that for every $\sigma\in\operatorname{Emb}(K/\mathbb{Q})$ with $\sigma\neq\tau$:
  • $i)$
    if $\sigma(\alpha_{2})\neq\alpha_{2}$, then $u,u^{\prime}\nmid N_{K}(\sigma(\alpha_{2})-\alpha_{2})$,
  • ii)ii)
    if $\sigma\neq\operatorname{Id}$, then $u,u^{\prime}\nmid\sigma(z)-z$,
  • $iii)$
    if $\tau(\alpha_{2,K})\neq\alpha_{2}$, then
    $u,u^{\prime}\nmid N_{K}\left(\left(\sigma(z)-z-\frac{\sigma(\alpha_{2,K})-\alpha_{2,K}}{\tau(\alpha_{2,K})-\alpha_{2,K}}\cdot(\tau(z)-z)\right)\right)$
Then, for any $s\in\mathbb{Z}/q_{1}\mathbb{Z}$, there exists $A_{s,u}\subset\{(r,t)\in\left(\mathbb{Z}/u\mathbb{Z}\right)\times\left(\mathbb{Z}/u\mathbb{Z}\right)\}$, and $A_{s,u^{\prime}}\subset\{(r,t)\in\left(\mathbb{Z}/u^{\prime}\mathbb{Z}\right)\times\left(\mathbb{Z}/u^{\prime}\mathbb{Z}\right)\}$, such that
  • $a)$
    for any $(r,t)$ such that $(r,t)\mod u\not\in A_{s,u}$, and $(r,t)\mod u\not\in A_{s,u^{\prime}}$, one has
    $\left|\sum_{x,y,w\text{ }(q_{1})}\left(\frac{F_{\operatorname{\mathbf{z}}}(X_{1},X_{2})}{q_{1}}\right)\left(\frac{F_{\operatorname{\mathbf{z}}}(w,y+r)}{q_{1}}\right)e_{q_{1}}(s(x-w)+ty)\right|\ll_{n}q_{1}^{3/2}(s,t,q_{1})^{\frac{1}{2}}.$
  • $b)$
    $A_{s,u}=O_{n}(1)$, and $A_{s,u^{\prime}}=O_{n}(1)$.
We are ready to bound the contribution of $\Sigma_{2B}$. First we recall that $q_{1}=uu^{\prime}$ with $u,u^{\prime}\in\mathcal{U}$. From now one we set
$\mathcal{U}=\left\{u:\begin{matrix}U\leq u\leq 2U\\ u\nmid N_{K}(\sigma(\alpha_{2})-\alpha_{2}),\text{ }\forall\sigma:\sigma(\alpha_{2})\neq\alpha_{2}\\ u\nmid N_{K}(\sigma(z)-z),\text{ }\forall\sigma\in\operatorname{Emb}(K/\mathbb{Q})\setminus\{\operatorname{Id}\}\\ u\nmid N_{K}\left(\left(\sigma(z)-z-\frac{\sigma(\alpha_{2,K})-\alpha_{2,K}}{\tau(\alpha_{2,K})-\alpha_{2,K}}\cdot(\tau(z)-z)\right)\right),\text{ }\forall\sigma,\tau\in\operatorname{Emb}(K/\mathbb{Q}):\tau(\alpha_{2,K})\neq\alpha_{2}\end{matrix}\right\},$
Moreover, we assume that $H\leq U$: in this way, it is guaranteed that $h\neq 0$, implies that $h\neq 0\text{ }(u)$ for every $U\leq u$. Since the bound for $W(u_{1}\overline{q_{2}},u_{2},q_{2}h;q_{1})$ in Corollary 4.13 depends on the G.C.D. of $q_{1},u_{1}\overline{q_{2}},$ and $u_{2}\overline{q_{2}}$. We have to handle two cases
  • $i)$
    the case when $(u_{1},q_{1})=1$. We start by applying Corollary 4.13, getting
    $\begin{split}\Sigma_{2B}&\ll\frac{1}{q_{1}}\sum_{\begin{subarray}{c}h\in[-H,H]\\ h\neq 0\end{subarray}}\left|\sum_{\begin{subarray}{c}u_{2}\in\mathbb{Z}\\ (q_{2}h,u_{2})\not\in A_{u_{1}\overline{q_{2}},u}\\ (q_{2}h,u_{2})\not\in A_{u_{1}\overline{q_{2}},u^{\prime}}\end{subarray}}\hat{g}_{h}\left(\frac{u_{2}}{q_{1}}\right)W(u_{1}\overline{q_{2}},u_{2},q_{2}h;q_{1})\right|\\ &\quad+\frac{1}{q_{1}}\sum_{\begin{subarray}{c}h\in[-H,H]\\ h\neq 0\end{subarray}}\left|\sum_{\begin{subarray}{c}u_{2}\in\mathbb{Z}\\ (q_{2}h,u_{2})\in A_{u_{1}\overline{q_{2}},u}\end{subarray}}\hat{g}_{h}\left(\frac{u_{2}}{q_{1}}\right)W(u_{1}\overline{q_{2}},u_{2},q_{2}h;q_{1})\right|\\ &\quad+\frac{1}{q_{1}}\sum_{\begin{subarray}{c}h\in[-H,H]\\ h\neq 0\end{subarray}}\left|\sum_{\begin{subarray}{c}u_{2}\in\mathbb{Z}\\ (q_{2}h,u_{2})\in A_{u_{1}\overline{q_{2}},u^{\prime}}\end{subarray}}\hat{g}_{h}\left(\frac{u_{2}}{q_{1}}\right)W(u_{1}\overline{q_{2}},u_{2},q_{2}h;q_{1})\right|\\ &\ll_{n}Hq_{1}^{\frac{3}{2}}+q_{1}^{2}\cdot\frac{1}{q_{1}}\sum_{\begin{subarray}{c}h\in[-H,H]\\ h\neq 0\end{subarray}}\sum_{\begin{subarray}{c}u_{2}\in\mathbb{Z}\\ (q_{2}h,u_{2})\in A_{u_{1}\overline{q_{2}},u}\end{subarray}}\left|\hat{g}_{h}\left(\frac{u_{2}}{q_{1}}\right)\right|\\ &\quad+q_{1}^{2}\cdot\frac{1}{q_{1}}\sum_{\begin{subarray}{c}h\in[-H,H]\\ h\neq 0\end{subarray}}\sum_{\begin{subarray}{c}u_{2}\in\mathbb{Z}\\ (q_{2}h,u_{2})\in A_{u_{1}\overline{q_{2}},u^{\prime}}\end{subarray}}\left|\hat{g}_{h}\left(\frac{u_{2}}{q_{1}}\right)\right|,\end{split}$
    Provided that we have $q_{1}\geq D_{K}^{\frac{1}{n}-\delta_{2,K}}$; now we have
    $\begin{split}\sum_{\begin{subarray}{c}h\in[-H,H]\\ h\neq 0\end{subarray}}\sum_{\begin{subarray}{c}u_{2}\in\mathbb{Z}\\ (q_{2}h,u_{2})\in A_{u_{1}\overline{q_{2}},u}\end{subarray}}\left|\hat{g}_{h}\left(\frac{u_{2}}{q_{1}}\right)\right|&\leq\sum_{(r,t)\in A_{u_{1}\overline{q_{2}},u}}\sum_{\begin{subarray}{c}h\in[-H,H]\\ h\neq 0\\ q_{2}h=r\text{ }(u)\end{subarray}}\sum_{\begin{subarray}{c}u_{2}\in\mathbb{Z}\\ u_{2}=t\text{ }(u)\end{subarray}}\left|\hat{g}_{h}\left(\frac{u_{2}}{q_{1}}\right)\right|\\ &\ll D_{K}^{\frac{1}{n}-\delta_{2,K}}\sum_{(r,t)\in A_{u_{1}\overline{q_{2}},u}}\sum_{\begin{subarray}{c}h\in[-H,H]\\ q_{2}h=r\text{ }(u)\end{subarray}}\sum_{\begin{subarray}{c}u_{2}\in\mathbb{Z}\\ u_{2}=t\text{ }(u)\end{subarray}}\left|1+\frac{u_{2}D_{K}^{\frac{1}{n}-\delta_{2,K}}}{q_{1}}\right|^{-2}\\ &\ll_{n}D_{K}^{\frac{1}{n}-\delta_{2,K}},\end{split}$
    provided that $\frac{q_{1}}{u}\leq D_{K}^{\frac{1}{n}-\delta_{2,K}}$, i.e. $q_{1}^{\frac{1}{2}}\ll D_{K}^{\frac{1}{n}-\delta_{2,K}}$. Moreover, recalling that $H=\lfloor\frac{D_{K}^{\frac{1}{n}-\delta_{2,K}}}{q_{2}}\rfloor$, we conclude that
    $\begin{split}q_{1}^{2}\cdot\frac{1}{q_{1}}\sum_{\begin{subarray}{c}h\in[-H,H]\end{subarray}}\sum_{\begin{subarray}{c}u_{2}\in\mathbb{Z}\\ (q_{2}h,u_{2})\in A_{u_{1}\overline{q_{2}},u}\end{subarray}}\left|\hat{g}_{h}\left(\frac{u_{2}}{q_{1}}\right)\right|&\ll_{n}D_{K}^{\frac{1}{n}-\delta_{2,K}}q_{1}\\ &\leq Hq_{2}q_{1}\\ &\leq Hq_{1}^{\frac{3}{2}},\end{split}$
    provided that $\max\{q_{2}^{2},D_{K}^{\frac{1}{n}-\delta_{2,K}}\}\leq q_{1}\leq D_{K}^{\frac{2}{n}-2\delta_{2,K}}$ (notice that $\max\{q_{2}^{2},D_{K}^{\frac{1}{n}-\delta_{2,K}}\}\leq q_{1}$ gurantees that $H\leq q_{1}$). Hence, in this case, we have
    $\Sigma_{2B}\ll_{n}Hq_{1}^{\frac{3}{2}},$
    provided that $\max\{q_{2}^{2},D_{K}^{\frac{1}{n}-\delta_{2,K}}\}\leq q_{1}\leq D_{K}^{\frac{2}{n}-2\delta_{2,K}}$.
  • ii)ii)
    The case $(u_{1},q_{1})>1$. We start by writing
    $\begin{split}\Sigma_{2B}&\ll\frac{1}{q_{1}}\sum_{\begin{subarray}{c}h\in[-H,H]\\ h\neq 0\end{subarray}}\left|\sum_{\begin{subarray}{c}u_{2}\in\mathbb{Z}\\ (u_{2},q_{1})=1\end{subarray}}\hat{g}_{h}\left(\frac{u_{2}}{q_{1}}\right)W(u_{1}\overline{q_{2}},u_{2},q_{2}h;q_{1})\right|\\ &\quad+\frac{1}{q_{1}}\sum_{\begin{subarray}{c}h\in[-H,H]\\ h\neq 0\end{subarray}}\left|\sum_{\begin{subarray}{c}u_{2}\in\mathbb{Z}\\ (u_{2},q_{1})>1\end{subarray}}\hat{g}_{h}\left(\frac{u_{2}}{q_{1}}\right)W(u_{1}\overline{q_{2}},u_{2},q_{2}h;q_{1})\right|\end{split}$ (17)
    observe that the contribution of the $u_{2}$ coprime $q_{1}$ can be bounded as in $(i)$, getting
    $\frac{1}{q_{1}}\sum_{\begin{subarray}{c}h\in[-H,H]\\ h\neq 0\end{subarray}}\left|\sum_{\begin{subarray}{c}u_{2}\in\mathbb{Z}\\ (u_{2},q_{1})=1\end{subarray}}\hat{g}_{h}\left(\frac{u_{2}}{q_{1}}\right)W(u_{1}\overline{q_{2}},u_{2},q_{2}h;q_{1})\right|\ll_{n}Hq_{1}^{\frac{3}{2}},$
[…]
    $\begin{split}\frac{1}{q_{1}}&\sum_{\begin{subarray}{c}h\in[-H,H]\\ h\neq 0\end{subarray}}\left|\sum_{\begin{subarray}{c}u_{2}\in\mathbb{Z}\\ (u_{2},q_{1})>1\end{subarray}}\hat{g}_{h}\left(\frac{u_{2}}{q_{1}}\right)W(u_{1}\overline{q_{2}},u_{2},q_{2}h;q_{1})\right|\\ &\ll_{n}q_{1}\sum_{\begin{subarray}{c}h\in[-H,H]\\ h\neq 0\end{subarray}}\sum_{\begin{subarray}{c}u_{2}\in\mathbb{Z}\\ (u_{2},q_{1})>1\end{subarray}}\left|\hat{g}_{h}\left(\frac{u_{2}}{q_{1}}\right)\right|\\ &\ll H\cdot q_{1}\cdot D_{K}^{\frac{1}{n}-\delta_{2,K}},\end{split}$
    provided that $q_{1}^{\frac{1}{2}}\leq D_{K}^{\frac{1}{n}-\delta_{2,K}}$. Thus
    $\begin{split}\frac{1}{q_{1}}\sum_{\begin{subarray}{c}h\in[-H,H]\end{subarray}}\left|\sum_{\begin{subarray}{c}u_{2}\in\mathbb{Z}\\ (u_{2},q_{1})>1\end{subarray}}\hat{g}_{h}\left(\frac{u_{2}}{q_{1}}\right)W(u_{1}\overline{q_{2}},u_{2},q_{2}h;q_{1})\right|&\ll_{n}HD_{K}^{\frac{1}{n}-\delta_{2,K}}q_{1},\end{split}$
    provided that $\max\{q_{2}^{2},D_{K}^{\frac{1}{n}-\delta_{2,K}}\}\leq q_{1}\leq D_{K}^{\frac{2}{n}-2\delta_{2,K}}$.
Combining the estimates for $\Sigma_{1},\Sigma_{2A}$, and $\Sigma_{2B}$, we conclude
Lemma 4.14.
Let $\operatorname{\mathbf{z}}\in\mathbb{Z}^{n-2}\setminus\{\boldsymbol{0}\}$ satisfying the following condition:
  • $i)$
    For every $\sigma,\tau\in\operatorname{Emb}(K/\mathbb{Q})$, with $\sigma\neq\tau$ and $\tau(\alpha_{2,K})\neq\alpha_{2,K}$,
    $\sigma(z)-z\neq\frac{\sigma(\alpha_{2,K})-\alpha_{2,K}}{\tau(\alpha_{2,K})-\alpha_{2,K}}(\tau(z)-z),$
  • ii)ii)
    For every $\sigma\in\operatorname{Emb}(K/\mathbb{Q})\setminus\{\operatorname{Id}\}$, $z\not\in K^{\sigma}$.
Then, provided that $H<q_{1}$ and $q_{2}^{2}=q_{1}$, one has
$S(\operatorname{\mathbf{z}},u_{1};q_{1}q_{2})\ll_{n}\begin{cases}D_{K}^{\frac{1}{2n}-\frac{\delta_{2,K}}{2}}q_{1}&\text{ if }(u_{1},q_{1})=1\\ \frac{D_{K}^{\frac{1}{n}-\delta_{2,K}}q_{1}}{q_{2}^{\frac{1}{2}}}&\text{otherwise.}\par\end{cases}$
provided that $\max\{q_{2}^{2},D_{K}^{\frac{1}{n}-\delta_{2,K}}\}\leq q_{1}\leq D_{K}^{\frac{2}{n}-2\delta_{2,K}}$, and $D^{\frac{1}{2n}-\delta_{2,K}}\leq q_{2}q_{1}^{\frac{1}{2}}$
Then we can finally bound the main sieve. From now on we set
[…]
Moreover, we require $D_{K}^{\frac{1}{2n}-\frac{\delta_{2,K}}{2}}\leq U\leq D_{K}^{\frac{1}{n}-\delta_{2,K}}$ and $U=V^{2}$, and $UV^{2}\leq D_{K}^{\frac{1}{n}}\leq(UV)^{2}$. We have
$S(\operatorname{\mathbf{z}};q_{1}q_{2})=\frac{1}{q_{1}q_{2}}\sum_{\begin{subarray}{c}u_{1}\in\mathbb{Z}\\ (u_{1},q_{1})=1\end{subarray}}\hat{g}_{1}\left(\frac{u_{1}}{q_{1}q_{2}}\right)S(\operatorname{\mathbf{z}},u_{1};q_{1}q_{2})+\frac{1}{q_{1}q_{2}}\sum_{\begin{subarray}{c}u_{1}\in\mathbb{Z}\\ (u_{1},q_{1})>1\end{subarray}}\hat{g}_{1}\left(\frac{u_{1}}{q_{1}q_{2}}\right)S(\operatorname{\mathbf{z}},u_{1};q_{1}q_{2})$
We start by dealing with the contribution of those $u_{1}$ such that $(u_{1},q_{1})>1$. An application of Lemma 4.14 combined with our assumption that $UV^{2}\leq D_{K}^{\frac{1}{n}}$, gives
$\begin{split}\frac{1}{q_{1}q_{2}}\sum_{\begin{subarray}{c}u_{1}\in\mathbb{Z}\\ (u_{1},q_{1})>1\end{subarray}}\hat{g}_{1}\left(\frac{u_{1}}{q_{1}q_{2}}\right)S(\operatorname{\mathbf{z}},u_{1};q_{1}q_{2})&\ll_{n}\frac{D^{\frac{2}{n}-\delta_{2,K}}}{q_{2}^{\frac{3}{2}}}\sum_{\begin{subarray}{c}u_{1}\in\mathbb{Z}\\ (u_{1},q_{1})>1\end{subarray}}\left|\hat{g}_{1}\left(\frac{u_{1}}{q_{1}q_{2}}\right)\right|\\ &\ll\frac{D^{\frac{2}{n}-\delta_{2,K}}}{UV}\end{split}$
Let us turn our attention those $u_{1}$ such that $(u_{1},q_{1})=1$. Applying Lemma 4.14, combined with our assumption that $(UV)^{2}\geq D_{K}^{\frac{1}{n}}$, gives
$\begin{split}\frac{1}{q_{1}q_{2}}\sum_{\begin{subarray}{c}u_{1}\in\mathbb{Z}\\ (u_{1},q_{1})=1\end{subarray}}\hat{g}_{1}\left(\frac{u_{1}}{q_{1}q_{2}}\right)S(\operatorname{\mathbf{z}},u_{1};q_{1}q_{2})&\ll_{n}D_{K}^{\frac{1}{2n}-\frac{\delta_{2,K}}{2}}U^{2}\end{split}$
Combining the two bounds, we get
$S(\operatorname{\mathbf{z}};q_{1}q_{2})\ll_{n}D_{K}^{\frac{1}{2n}-\frac{\delta_{2,K}}{2}}U^{2}+\frac{D^{\frac{2}{n}-\delta_{2,K}}}{UV}$
[…]
$\frac{1}{A^{2}}\sum_{v,v^{\prime}\in\mathcal{V}}\sum_{\begin{subarray}{c}u,u^{\prime}\mathcal{U}\\ u\neq u^{\prime}\end{subarray}}\left|\sum_{n\in\mathbb{Z}}\left(\frac{n}{uu^{\prime}}\right)\left(\frac{n}{uu^{\prime}}\right)\omega_{\operatorname{\mathbf{z}}}(n)\right|\ll_{n}D_{K}^{\frac{1}{2n}-\frac{\delta_{2,K}}{2}}U^{2}+\frac{D^{\frac{2}{n}-\delta_{2,K}}}{UV}.$ (18)
4.4.1. Proof of Lemma 4.12: the prime sieve
We briefly discuss now how to handle the Prime sieve
$N=\frac{U}{A^{2}}\sum_{\begin{subarray}{c}v,v^{\prime}\mathcal{V}\\ v\neq v^{\prime}\end{subarray}}\left|\sum_{n\in\mathbb{Z}}\left(\frac{n}{vv^{\prime}}\right)\omega_{\operatorname{\mathbf{z}}}(n)\right|.$
[…]
arguing as in Lemma 4.10, one gets
$\left|\sum_{n\in\mathbb{Z}}\left(\frac{n}{vv^{\prime}}\right)\omega_{\operatorname{\mathbf{z}}}(n)\right|\ll_{n}\frac{D_{K}^{\frac{2}{n}-\delta_{2,K}}}{vv^{\prime}}\sum_{\begin{subarray}{c}\operatorname{\mathbf{u}}\in\mathbb{Z}^{2}\end{subarray}}\left(1+\frac{D_{K}^{\frac{1}{n}}u_{1}}{vv^{\prime}}\right)^{-2}\cdot\left(1+\frac{D_{K}^{\frac{1}{n}-\delta_{2,K}}u_{2}}{vv^{\prime}}\right)^{-2}.$ (19)
[…]
$\left|\sum_{n\in\mathbb{Z}}\left(\frac{n}{vv^{\prime}}\right)\omega_{\operatorname{\mathbf{z}}}(n)\right|\ll_{n}\frac{D_{K}^{\frac{2}{n}-\delta_{2,K}}}{V^{2}}$
Hence,
$\frac{U}{A^{2}}\sum_{\begin{subarray}{c}v,v^{\prime}\mathcal{V}\\ v\neq v^{\prime}\end{subarray}}\left|\sum_{n\in\mathbb{Z}}\left(\frac{n}{vv^{\prime}}\right)\omega_{\operatorname{\mathbf{z}}}(n)\right|\ll_{n}\frac{D_{K}^{\frac{2}{n}-\delta_{2,K}}}{UV^{2}},$ (20)
which is acceptable.
4.4.2. Proof of Lemma 4.12: conclusion
[…]
$\sum_{k}\omega_{\operatorname{\mathbf{z}}}(k^{2})\ll_{n}\frac{D_{K}^{\frac{2}{n}-\delta_{2,K}}}{UV}+D_{K}^{\frac{1}{2n}-\frac{\delta_{2,K}}{2}}U^{2}.$
Imposing $U=V^{2}$, the bound above is optimal as soon as
$\frac{D_{K}^{\frac{2}{n}-\delta_{2,K}}}{V^{3}}=D_{K}^{\frac{1}{2n}-\frac{\delta_{2,K}}{2}}V^{4}$
which implies
$V=D_{K}^{\frac{1}{7n}+\frac{1}{14}\left(\frac{1}{n}-\delta_{2,K}\right)}$
and concludes the proof of Lemma 4.12.
4.4.3. Proof of Theorem 2.3
We can now prove Theorem 2.3. Again, thanks to Lemma 4.1, we may assume that $K$ does not contain a subfield of index 22. We recall that
$\mathcal{C}=\left[-D_{K}^{\frac{1}{n}-\delta_{3,K}},D_{K}^{\frac{1}{n}-\delta_{3,K}}\right]\times\cdots\times\left[-D_{K}^{\frac{1}{n}-\delta_{n,K}},D_{K}^{\frac{1}{n}-\delta_{n,K}}\right],$
and that
[…]
We start by writing
$\begin{split}h_{2}(K)&\ll\#\left\{(y,z_{1},...,z_{n})\in\mathbb{Z}^{n+1}:\begin{matrix}|z_{i}|\ll D_{K}^{\frac{1}{n}-\delta_{i,K}},\quad\text{for }i=1,...,n\\ y^{2}=N_{K}(z_{1}+z_{2}\alpha_{2,K}+\cdots+z_{n}\alpha_{n,K})\end{matrix}\right\}\\ &=\sum_{\operatorname{\mathbf{z}}\in\mathcal{C}\cap\mathbb{Z}^{n-2}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1\\ &=\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\cap Z_{\square}\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1+\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\\ \operatorname{\mathbf{z}}\in Z_{F}\setminus Z_{\square}\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1+\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\cap\mathbb{Z}^{n-2}\\ \operatorname{\mathbf{z}}\not\in Z_{\square}\cup Z_{F}\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1.\end{split}$ (21)
For the first two terms we have the bounds
$\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\cap Z_{\square}\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1=\begin{cases}0&\text{if $n=5,6$}\\ O_{n,\varepsilon}\left(D_{K}^{\frac{1}{4}+\frac{1}{n}}\right)&\text{if $n>6$}\end{cases},\qquad\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\\ \operatorname{\mathbf{z}}\in Z_{F}\setminus Z_{\square}\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1\ll_{n,\varepsilon}D_{K}^{\frac{1}{2}-\frac{1}{2n}-\frac{1}{7n}},$
[…]
$\begin{split}\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\cap\mathbb{Z}^{n-2}\\ \operatorname{\mathbf{z}}\not\in Z_{\square}\cup Z_{F}\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1&\ll_{n,\varepsilon}\#(\mathcal{C}\cap\mathbb{Z}^{n-2})\cdot D_{K}^{\frac{45}{28n}-\frac{11}{14}\delta_{2,K}}\\ &\ll D_{K}^{\sum_{j=3}^{n}\left(\frac{1}{n}-\delta_{j,K}\right)+\frac{1}{2}-\frac{1}{2n}-\left(\frac{1}{7n}-\frac{3\delta_{2,K}}{14}\right)+\varepsilon}\\ &=D_{K}^{\frac{1}{2}-\frac{1}{2n}-\left(\frac{1}{7n}-\frac{3\delta_{2,K}}{14}\right)+\varepsilon}.\end{split}$
which completes the proof of Theorem 2.3.
5. The Relevant exponential sums.
This section is dedicated to prove the key exponential sums used in proof of Lemma 4.9 and Proposition 2.4.
5.1. Notions of the formalism of $\ell$-adic trace functions
We first recall some notions of the formalism of $\ell$-adic trace functions. We are assuming the definition of a constructible $\ell$-adic sheaf, for which, as well as for a general introduction on the subject, we refer the reader to [Del80, Section $1.1$] and [Kat80, Section $3.4.2$]. We start by introducing the notion of a trace function attached to a constructible $\ell$-adic sheaf as in [Kat90, $7.3.7$].
In the following $p,\ell>2$ and $q=p^{k}$ for some $k\geq 1$ are distinct prime numbers and $\iota:\overline{\mathbb{Q}}_{\ell}\hookrightarrow\mathbb{C}$ is a fixed embedding. Let $\mathcal{F}$ be a constructible $\ell$-adic sheaf on $\overline{\mathbb{A}}_{\mathbb{F}_{q}}^{1}$. For any $x\in\overline{\mathbb{A}}_{\mathbb{F}_{q}}^{1}(\mathbb{F}_{q^{n}})$ one defines
$t_{\mathcal{F},n}(x):=\iota(\operatorname{Tr}(\operatorname{Fr}_{q^{n}}|\mathcal{F}_{\overline{x}})),$
where $\operatorname{Fr}_{q^{n}}$ is the geometric Frobenius automorphism of $\mathbb{F}_{q^{n}}$ and $\mathcal{F}_{\overline{x}}$ is the stalk of $\mathcal{F}$ at a geometric point $\overline{x}$ over $x$. The function $t_{\mathcal{F},n}$ is called the trace function attached to $\mathcal{F}$ over $\mathbb{F}_{q^{n}}^{1}$. If there is not ambiguity, we denote by $t_{\mathcal{F}}$ the trace function $t_{\mathcal{F},1}$.
Definition 5.1.
Let $\mathcal{F}$ be a constructible $\ell$-adic sheaf on $\overline{\mathbb{A}}_{\mathbb{F}_{q}}^{1}$ and $j:U\hookrightarrow\overline{\mathbb{A}}_{\mathbb{F}_{q}}^{1}$ the largest dense open subset of $\overline{\mathbb{A}}_{\mathbb{F}_{q}}^{1}$ where $\mathcal{F}$ is lisse.
  • $i)$
    The sheaf $\mathcal{F}$ is said to be a middle-extension $\ell$-adic sheaf if $\mathcal{F}=j_{*}j^{*}\mathcal{F}$ (see [Kat80, $4.4$, $4.5$] for the definition of $j_{*}j^{*}\mathcal{F}$ and its basic properties).
  • ii)ii)
    The sheaf $\mathcal{F}$ is said to be a middle-extension $\ell$-adic sheaf, punctually pure of weight 0 if it is a middle-extension sheaf and if for every $n\geq 1$ and every $x\in U(\mathbb{F}_{q^{n}})$, the images of the eigenvalues of $(\operatorname{Fr}_{q^{n}}|\mathcal{F}_{\overline{x}})$ via the fixed embedding $\iota:\overline{\mathbb{Q}}_{\ell}\hookrightarrow\mathbb{C}$, are complex numbers of modulus 11.
Remark 1.
By [FKM15, Definition $1.12$], we will refer to trace functions attached to punctually pure of weight 0 middle-extension $\ell$-adic sheaves on $\overline{\mathbb{A}}_{\mathbb{F}_{q}^{1}}$ as trace functions.
Definition 5.2 ([FKM19], pp. 464-6).
Let $\mathcal{F}$ be a constructible $\ell$-adic sheaf on $\overline{\mathbb{A}}_{\mathbb{F}_{q}^{1}}^{1}$ and $j:U\hookrightarrow\overline{\mathbb{A}}_{\mathbb{F}_{q}^{1}}^{1}$ the largest dense open subset of $\overline{\mathbb{A}}_{\mathbb{F}_{q}^{1}}^{1}$ where $\mathcal{F}$ is lisse. The conductor of $\mathcal{F}$ is defined as
$c(\mathcal{F}):=\operatorname{rank}(\mathcal{F})+|\operatorname{sing}(\mathcal{F})|+\sum_{x\in\overline{{\mathbb{P}}}_{\mathbb{F}_{q}^{1}}^{1}(\overline{\mathbb{F}}_{q})}\text{Swan}_{x}(j_{*}j^{*}\mathcal{F})+\dim H_{c}^{0}(\overline{\mathbb{A}}_{\mathbb{F}_{q}}^{1},\mathcal{F}),$
where
  1. $i)$
    $\operatorname{rank}(\mathcal{F}):=\dim\mathcal{F}_{x}$, for any $x$ where $\mathcal{F}$ is lisse.
  2. ii)ii)
    $\operatorname{sing}(\mathcal{F}):=\{x\in\overline{\mathbb{P}}_{\mathbb{F}_{q}}^{1}(\overline{\mathbb{F}}_{q}):\mathcal{F}\text{ is not lisse at }x\}$.
  3. $iii)$
    For any $x\in\overline{\mathbb{P}}_{\mathbb{F}_{q}^{1}}^{1}(\overline{\mathbb{F}}_{q})$, $\operatorname{Swan}_{x}(j_{*}j^{*}\mathcal{F})$ is the Swan conductor of $\mathcal{F}$ at $x$ (see [Kat88, Chapter 11] for the definition of the Swan conductor).
Remark 2.
Recall that if $\mathcal{F}$ is a middle-extension sheaf then
$c(\mathcal{F})=\operatorname{rank}(\mathcal{F})+|\operatorname{sing}(\mathcal{F})|+\sum_{x\in\overline{{\mathbb{P}}}_{\mathbb{F}_{q}}^{1}(\overline{\mathbb{F}}_{q})}\text{Swan}_{x}(\mathcal{F}),$
since in this case $\mathcal{F}\cong j_{*}j^{*}\mathcal{F}$ and $\dim H_{c}^{0}(\overline{\mathbb{A}}_{\mathbb{F}_{q}}^{1},\mathcal{F})=0$.
Finally, we recall the following construction:
Example 5.3.
let $F[X,Y]\in\mathbb{F}_{q}[X,Y]$ be a square-free polynomial of degree $n$. Then for any non trivial multiplicative character over $\mathbb{F}_{q}^{\times}$, $\chi$, one can consider the following $t:\mathbb{F}_{q}\rightarrow\mathbb{C}$
$x\mapsto\sum_{x\in\mathbb{F}_{q}}\chi(F(x,y)).$
It turns out that we can see this function as a trace function: combining [FKM19, Definition $5.9$, Theorem $5.10$] and arguing like in [FKM19, Corollary $5.10$] one shows that there exists a middle extension $\ell$-adic sheaf $\mathcal{H}_{F,\chi}$ pure of weight 11 such that
$t_{\mathcal{H}_{F,\chi},k}(x)=\sum_{y\in\mathbb{F}_{q^{k}}}\chi_{k}(F(x,y))+O(1)\quad\text{ for all }x\in\mathbb{F}_{q^{k}}\setminus(\operatorname{sing}(\mathcal{H}_{F,\chi})\cap\mathbb{F}_{q^{k}}),$
where $\chi_{k}=\chi\circ\text{Norm}_{\mathbb{F}_{q^{k}}/\mathbb{F}_{q}}$. Moreover, one has that
  • $i)$
    $\operatorname{sing}(\mathcal{H}_{F,\chi})\cap\overline{\mathbb{A}}_{\mathbb{F}_{q}}^{1}(\overline{\mathbb{F}}_{q})=\{y\in\overline{\mathbb{A}}_{\mathbb{F}_{q}}^{1}(\overline{\mathbb{F}}_{q}):F(X,y)\text{ is not square-free}\}$
  • ii)ii)
    $\operatorname{rank}(\mathcal{H}_{F,\chi})=\dim(H^{1}(\overline{\mathbb{A}}_{\mathbb{F}_{q}}^{1},\mathcal{L}_{\chi(F(x,Y))}))$ for any $x$ where $\mathcal{H}$ is lisse.
  • $iii)$
    $C(\mathcal{H}_{F,\chi})\ll_{n}1$
5.1.1. A 11-dimensional exponential sum
We conclude this section by recording the following Lemma about one dimensional exponential sums
Lemma 5.4.
Let $F[X,Y]\in\mathbb{F}_{q}[X,Y]$ be a square free polynomial. Then
$\left|\sum_{x\in\mathbb{F}_{p}}\left(\frac{F(x,y)}{p}\right)e_{p}(sx)\right|\ll_{\deg F}p^{\frac{1}{2}},$
for all but $O_{\deg F}(1)$ values of $y$.
Proof.
If $s\neq 0$, then the result follows from [We48]. If $s=0$, then from [IK04, Theorem $11.23$], we have that
$\left|\sum_{x\in\mathbb{F}_{p}}\left(\frac{F(x,y)}{p}\right)\right|\ll_{\deg F}p^{\frac{1}{2}},$
for every $y$ such that $F(X,y)\in\mathbb{F}_{p}[X]$ is not a square. it rests to show that $F(X,y)$ can be a square only for $O_{\deg F}(1)$ values of $y$: since we are assuming $F(X,Y)\in\mathbb{F}_{q}(Y)[X]$ is square free it follows that $\text{disc}(F)\in\mathbb{F}[Y]\neq 0$. Moreover,
$F(X,y)\text{ is square-free}\Longleftrightarrow\text{disc}(F)(y)\neq 0.$
Thus $F(X,y)$ is square-free for but $O_{\deg F}(1)$ values of $y$. ∎
5.2. Hooley’s Theorem for exponential sums.
Both the proofs of Proposition 2.4 and Lemma 4.9 rely on a general procedure due to Hooley [Ho82] (see also ([BB23, Theorem A.1])), which allows one to estimate a very general family of exponential sums over a finite field, provided that one can control the seconds moment of an appropriate counting function.
Theorem 5.5 (Hooley).
Let $F,G_{1},\dots,G_{k}\in\mathbb{Z}[X_{1},\dots,X_{m}]$ be polynomials of degree at most $d$ and let
$S=\sum_{\begin{subarray}{c}x\in\mathbb{F}_{p}^{m}\\ G_{1}(\mathbf{x})=\cdots=G_{k}(\mathbf{x})=0\end{subarray}}e_{p}(F(\mathbf{x})).$
For each $r\geq 1$ and $\tau\in\mathbb{F}_{p^{r}}$, write
$N_{r}(\tau)=\#\left\{\mathbf{x}\in\mathbb{F}_{p^{r}}^{m}:G_{1}(\mathbf{x})=\dots=G_{k}(\mathbf{x})=0,F(\mathbf{x})=\tau\right\}.$ (22)
If there exist $N_{r}\in\mathbb{R}$ such that
$\sum_{\tau\in\mathbb{F}_{p^{r}}}|N_{r}(\tau)-N_{r}|^{2}\ll_{d,k,m}p^{\kappa r},$
where $\kappa\in\mathbb{Z}$ is independent of $r$, then $S\ll_{d,k,m}p^{\kappa/2}$.
5.3. Proof of Lemma 4.9.
We are now ready to prove Lemma 4.9 which is a direct consequence of:
Lemma 5.6.
Let $F[X,Y]\in\mathbb{F}_{q}[X,Y]$ be a square-free polynomial such that
$F[X,Y]=\prod_{i=1}^{n}(X+a_{i}Y+b_{i})\text{ over }\overline{\mathbb{F}}_{q}.$
Moreover assume that $a_{i}\neq a_{1}$ for some $i\in\{2,...,n\}$. Then for any $s,t\in\mathbb{F}_{p}$, we have
$\left|\sum_{x,y\in\mathbb{F}_{p}}\left(\frac{F(x,y)}{p}\right)e_{p}(sx+ty)\right|\ll_{n}p(p,r,s)^{\frac{1}{2}}.$
Proof.
We start by dealing with the case when $r=s=0$: we have
$\sum_{x,y\in\mathbb{F}_{p}}\left(\frac{F(x,y)}{p}\right)=\sum_{y\in\mathbb{F}_{p}}\sum_{x\in\mathbb{F}_{p}}\left(\frac{F(x,y)}{p}\right).$
Thanks to Lemma 5.4, we get
$\left|\sum_{x\in\mathbb{F}_{p}}\left(\frac{F(x,y)}{p}\right)\right|\ll_{n}\sqrt{p},$ (23)
[…]
$\begin{split}\sum_{y\in\mathbb{F}_{p}}\sum_{x\in\mathbb{F}_{p}}\left(\frac{F(x,y)}{p}\right)&=\sum_{\begin{subarray}{c}y\in\mathbb{F}_{p}\setminus S\end{subarray}}\sum_{x\in\mathbb{F}_{p}}\left(\frac{F(x,y)}{p}\right)+\sum_{\begin{subarray}{c}y\in S\end{subarray}}\sum_{x\in\mathbb{F}_{p}}\left(\frac{F(x,y)}{p}\right)\\ &\ll_{n}p^{\frac{3}{2}}.\end{split}$
Now we deal with the case $(s,t)\neq(0,0)$. We may assume $s\neq 0$ (the case $t\neq 0$ is similar). We start by observing that
$\sum_{x,y\in\mathbb{F}_{p}}\left(\frac{F(x,y)}{p}\right)e_{p}(sx+ty)=\sum_{\begin{subarray}{c}w,x,y\in\mathbb{F}_{p}\\ w^{2}=F(x,y)\end{subarray}}e_{p}(sx+ty).$
With the goal of using Theorem 5.5 , we are going to compute
$\begin{split}N_{m}(k)&=\#\{(w,x,y)\in\mathbb{F}_{p^{m}}^{3}:w^{2}=F(x,y),\text{ }sx+ty=k\}|\\ &=\#\{(w,x)\in\mathbb{F}_{p^{m}}^{3}:w^{2}=F(x,s^{\prime}x+k^{\prime})\},\end{split}$
for every $m\geq 1$, where $s^{\prime}=-\overline{t}s$, and $k^{\prime}=\overline{t}k$. We claim that $F(X,s^{\prime}X+k^{\prime})$ is square-free for all but $O_{n}(1)$ values of $k^{\prime}$. With this in hand, let us see how we can finish the Lemma: if the polynomial $F(X,s^{\prime}X+k^{\prime})$ is square-free, then the polynomial $W^{2}=F(X,s^{\prime}X+k^{\prime})$ is irreducible. Provided that $F(X,s^{\prime}X+k^{\prime})$ is not a square, it follows from the Lang-Weil bound that
$\{(w,x)\in\mathbb{F}_{p^{m}}^{3}:w^{2}=F(x,s^{\prime}x+k^{\prime})\}=p^{m}+O_{n}(p^{\frac{m}{2}}).$
Thus,
$\sum_{k\in\mathbb{F}_{p^{m}}}|N_{m}(k)-p^{m}|^{2}\ll_{n}p^{2m},$
which leads to
$\left|\sum_{x,y\in\mathbb{F}_{p}}\left(\frac{F(x,y)}{p}\right)e_{p}(sx+ty)\right|\ll_{n}p,$
thanks to Theorem $5.$ It rests to prove that $F(X,s^{\prime}X+k^{\prime})$ is square-free for all but $O_{n}(1)$ values of $k^{\prime}$. First observe that
$F(X,s^{\prime}X+k^{\prime})=A_{s^{\prime},k^{\prime}}G_{s^{\prime},k^{\prime}}(X),$
where
$A_{s^{\prime},k^{\prime}}=\prod_{i:a_{i}s^{\prime}=-1}(b_{i}+a_{i}k^{\prime}),\quad G_{s^{\prime},k^{\prime}}(X)=\prod_{\begin{subarray}{c}i=1\\ a_{i}s^{\prime}\neq-1\end{subarray}}^{n}((1+a_{i}s^{\prime})X+a_{i}k^{\prime}+b_{i}).$
Since we are assuming that there exists $i$ such that $a_{i}\neq a_{1}$, then $G_{s^{\prime},k^{\prime}}(X)$ is a non-constant polynomial. Moreover, $F(X,s^{\prime}X+k^{\prime})$ is square-free if and only if $G_{s^{\prime},k^{\prime}}(X)$ is. On the other hand, $G_{s^{\prime},k^{\prime}}(X)$ is not square-free if and only if two distinct factors
$(1+a_{i}s^{\prime})X+a_{i}k^{\prime}+b_{i},\quad(1+a_{j}s^{\prime})X+a_{j}k^{\prime}+b_{j}$
are proportional. These two linear equations are proportional if and only if
$\frac{a_{i}k^{\prime}+b_{i}}{(1+a_{i}s^{\prime})}=\frac{a_{j}k^{\prime}+b_{j}}{(1+a_{j}s^{\prime})}$
which is true if and only if
$k^{\prime}(a_{i}-a_{j})=b_{j}(1+a_{i}s^{\prime})-b_{i}(1+a_{j}s^{\prime}).$
If $a_{i}-a_{j}\neq 0$, the equation above is satisfied for a unique value of $k^{\prime}$. If $a_{i}=a_{j}$, the equation above become $(b_{j}-b_{i})(1+a_{i}s^{\prime})=0$ which would implies $b_{j}-b_{i}=0$ since every factor appearing in the decomposition of $G_{s^{\prime},k^{\prime}}(X)$ satisfies $1+a_{i}s^{\prime}\neq 0$. On the other hand, since we are assuming that $F(X,Y)$ is square-free, if two factors
$X+a_{i}Y+b_{i},\quad X+a_{j}Y+b_{j},$
are such that $a_{i}=a_{j}$, then necessarily one has $b_{i}-b_{j}\neq 0$. Hence, $G_{s^{\prime},k^{\prime}}(X)$ is square free for all but $O_{n}(1)$ values of $k^{\prime}$ as claimed.
Proof of Lemma 4.9..
Since $F_{\operatorname{\mathbf{z}}}=N_{K}(X_{1}+X_{2}\alpha_{2}+z)$, with $z=z_{3}\alpha_{3}+\cdots+z_{n}\alpha_{n}$, it follows that
$F_{\operatorname{\mathbf{z}}}(X_{1},X_{2})=\prod_{\sigma\in\operatorname{Emb}(K/\mathbb{Q})}(X_{1}+X_{2}\sigma(\alpha_{2})+\sigma(z)).$
Now let $F_{\operatorname{\mathbf{z}}}=H_{\operatorname{\mathbf{z}}}^{2}G_{\operatorname{\mathbf{z}}}$ with $H_{\operatorname{\mathbf{z}}},G_{\operatorname{\mathbf{z}}}\in\mathbb{Q}[X,Y]$, monic and $G_{\operatorname{\mathbf{z}}}$ square-free. Since
$\sum_{x_{1},x_{2}\in\mathbb{F}_{p}}\left(\frac{F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})}{p}\right)e_{p}(sx_{1}+tx_{2})=\sum_{x_{1},x_{2}\in\mathbb{F}_{p}}\left(\frac{G_{\operatorname{\mathbf{z}}}(x_{1},x_{2})}{p}\right)e_{p}(sx_{1}+tx_{2})+O_{n}(p),$
it is enough to show that $G_{\operatorname{\mathbf{z}}}(X_{1},X_{2})$ satisfies the hypothesis of Lemma 5.6. Since we are assuming $F_{\operatorname{\mathbf{z}}}\mod p$ is not a square, it follows that
$G_{\operatorname{\mathbf{z}}}(X_{1},X_{2})=\prod_{\sigma\in\mathcal{E}}(X_{1}+X_{2}\sigma(\alpha_{2})+\sigma(z)),$
for some $\emptyset\neq\mathcal{E}\subset\operatorname{Emb}(K/\mathbb{Q})$. Let $\sigma\in\mathcal{E}$. Since $\sigma(\alpha_{2})\not\in\mathbb{Q}$, we can find $\tau\in\operatorname{Emb}(K/\mathbb{Q})$ such that $\tau(\sigma(\alpha_{2}))\neq\sigma(\alpha_{2})$. On the other hand,
$\tau((X_{1}+X_{2}\sigma(\alpha_{2})+\sigma(z))|\tau(G_{\operatorname{\mathbf{z}}})=G_{\operatorname{\mathbf{z}}}.$
Hence, $G_{\operatorname{\mathbf{z}}}$ satisfies the hypothesis of lemma 5.6, since it has two linear factors
$X_{1}+X_{2}\sigma(\alpha_{2})+\sigma(z),\quad X_{1}+X_{2}\tau\circ\sigma(\alpha_{2})+\tau\circ\sigma(z),$
such that $\sigma(\alpha_{2})\neq\tau\circ\sigma(\alpha_{2})$. ∎
5.4. Proof of Proposition 2.4
This section is devoted to bounding the exponential sum
$\sum_{(x,y,z)\in\mathbb{F}_{p}^{3}}\left(\frac{F(x,y)}{p}\right)\left(\frac{F(z,y+r)}{p}\right)e_{p}(s(x-z)+ty)$
for a given square-free polynomial $F(X,Y)$. We begin by establishing the following general estimate:
Lemma 5.7.
Let $F\in\mathbb{F}_{p}[X,Y]$ be a square-free polynomial with $\deg_{X}F\geq 1$. Then
$\left|\sum_{(x,y,z)\in\mathbb{F}_{p}^{3}}\left(\frac{F(x,y)}{p}\right)\left(\frac{F(z,y+r)}{p}\right)e_{p}(s(x-z)+ty)\right|\ll_{\deg F}p^{2}$
Proof.
We start by rewriting
$\begin{split}\sum_{(x,y,z)\in\mathbb{F}_{p}^{3}}&\left(\frac{F(x,y)}{p}\right)\left(\frac{F(z,y+r)}{p}\right)e_{p}(s(x-z)+ty)\\ &=\sum_{y\in\mathbb{F}_{p}}e_{p}(ty)\left(\sum_{x\in\mathbb{F}_{p}}\left(\frac{F(x,y)}{p}\right)e_{p}(sx)\right)\left(\sum_{z\in\mathbb{F}_{p}}\left(\frac{F(z,y+r)}{p}\right)e_{p}(-sz)\right).\end{split}$
Thanks to Lemma 5.4, there exists $S\subset\mathbb{F}_{p}$, of cardinality $S=O_{\deg F}(1)$ and such that
$\left|\sum_{x\in\mathbb{F}_{p}}\left(\frac{F(x,y)}{p}\right)e_{p}(sx)\right|\ll_{\deg F}\sqrt{p},\quad\left|\sum_{z\in\mathbb{F}_{p}}\left(\frac{F(z,y+r)}{p}\right)e_{p}(-sz)\right|\ll_{\deg F}\sqrt{p},$
for every $y\not\in S$. Then
$\begin{split}\sum_{(x,y,z)\in\mathbb{F}_{p}^{3}}&\left(\frac{F(x,y)}{p}\right)\left(\frac{F(z,y+r)}{p}\right)e_{p}(s(x-z)+ty)\\ &=\sum_{y\in\mathbb{F}_{p}\setminus S}e_{p}(ty)\left(\sum_{x\in\mathbb{F}_{p}}\left(\frac{F(x,y)}{p}\right)e_{p}(sx)\right)\left(\sum_{z\in\mathbb{F}_{p}}\left(\frac{F(z,y+r)}{p}\right)e_{p}(-sz)\right)\\ &\quad+\sum_{y\in S}e_{p}(ty)\left(\sum_{x\in\mathbb{F}_{p}}\left(\frac{F(x,y)}{p}\right)e_{p}(sx)\right)\left(\sum_{z\in\mathbb{F}_{p}}\left(\frac{F(z,y+r)}{p}\right)e_{p}(-sx)\right)\\ &\ll_{\deg F}p^{2}.\end{split}$
Lemma 5.8.
Let $F[X,Y]\in\mathbb{F}_{q}[X,Y]$ be a square-free polynomial such that
$F[X,Y]=\prod_{i=1}^{n}(X+a_{i}Y+b_{i})\text{ over }\overline{\mathbb{F}}_{q}.$
Moreover, assume that the following conditions hold:
  • $i)$
    There exists $i\in\{2,...,n\}$ such that $a_{1}\neq a_{i}$.
  • ii)ii)
    $\frac{a_{i}-a_{1}}{a_{j}-a_{1}}\cdot(b_{j}-b_{1})\neq b_{i}-b_{1}$ for any $i,j$ such that $a_{j}\neq a_{1}$ and $i\neq j$.
Then the sheaf $\mathcal{H}_{F,\chi}$ does not have any irreducible component that is geometrically trivial for every multiplicative character $\chi$ of $\mathbb{F}_{p}^{\times}$, i.e.
$\left|\sum_{(x,y)\in\mathbb{F}_{p}^{2}}\chi(F(x,y))\right|\ll_{n}p,$
for every multiplicative character $\chi$ of $\mathbb{F}_{p}^{\times}$.
Proof.
Let $\mathcal{H}_{F,\chi}=\bigoplus_{j}\mathcal{F}_{j}$ be the decomposition of $\mathcal{H}_{F,\chi}$ into irreducible components. Define $J=\{j:\mathcal{F}_{j}\cong_{\text{geom}}\mathbb{Q}_{z}\}$. For any $r\geq 1$, we have
$\sum_{x\in\mathbb{F}_{q^{r}}}t_{\mathcal{H}_{F,\chi},r}(x)=q^{\frac{3r}{2}}\sum_{j\in J}\omega_{j}^{r}+O(q^{r}),$
where $|\omega_{j}|=1$ for every $j\in J$, thanks to Deligne’s resolution of the Riemann hypothesis over finite fields ([FKMS19, Theorem $5.2$]). On the other hand, for every $m>0$, we have
$\begin{split}\frac{1}{R}\sum_{r\leq R}\left|\frac{1}{q^{\frac{3mr}{2}}}\sum_{x\in\mathbb{F}_{q^{mr}}}t_{\mathcal{H}_{F,\chi},mr}(x)\right|^{2}&=\frac{1}{R}\sum_{r\leq R}\left|\sum_{j\in J}w_{j}^{mr}\right|^{2}+\frac{1}{q^{\frac{mr}{2}}R}\end{split}$
Since $w_{j}$ are all root of unity we get
$\begin{split}\lim_{R\rightarrow\infty}\frac{1}{R}\sum_{r\leq R}\left|\sum_{j\in J}w_{j}^{mr}\right|^{2}\geq\#J.\end{split}$
which leads to
$\lim_{R\rightarrow\infty}\frac{1}{R}\sum_{r\leq R}\left|\frac{1}{q^{\frac{3mr}{2}}}\sum_{x\in\mathbb{F}_{q^{mr}}}t_{\mathcal{H}_{F,\chi},mr}(x)\right|^{2}\geq\#J.$ (24)
We claim that we can find $m$ such that for every $r\geq 1$,
$\sum_{x\in\mathbb{F}_{q^{mr}}}t_{\mathcal{H}_{F,\chi},mr}(x)=O_{n}(q^{mr}).$
[…]
$t_{\mathcal{H}_{F,\chi},m}(x)=\sum_{y\in\mathbb{F}_{q^{m}}}\chi_{m}(F(x,y))+O(1)\quad\text{ for all }x\in\mathbb{F}_{q^{m}}\setminus(\mathrm{Sing}(\mathcal{H}_{F,\chi})\cap\mathbb{F}_{q^{m}}),$
So it is enough to prove that for a suitable $m\geq 1$ and for every $r\geq 1$,
$\sum_{x,y\in\mathbb{F}_{q^{mr}}}\chi_{mr}(F(x,y))=O_{n}(q^{mr}).$
Let $m$ be an integer such that $F[X,Y]$ totally splits over $\mathbb{F}_{q^{m}}$. Then for every $r\geq 1$ we have
$F(X,Y)=\prod_{i=1}^{n}(X+a_{i}Y+b_{i}),\text{ over $\mathbb{F}_{q^{mr}}$}.$
Hence
$\sum_{x,y\in\mathbb{F}_{q^{mr}}}\chi_{mr}(F(x,y))=\sum_{x,y\in\mathbb{F}_{q^{mr}}}\prod_{i}^{n}\chi_{mr}(x+a_{i}y+b_{i}).$
Using the change of variables $z=x+a_{1}y$, we can rewrite
$F(x,y)=g(z)G(z,y),$
where
$g(z)=\prod_{i:a_{i}=a_{1}}(z+b_{i}),\quad G(z,y)=\prod_{i:a_{i}\neq a_{1},}(z+(a_{i}-a_{1})y+b_{i}).$
Since we are assuming that
  • $\bullet$
    the polynomial
    $F[X,Y]=\prod_{i=1}^{n}(X+a_{i}Y+b_{i})\text{ over }\overline{\mathbb{F}}_{q},$
    is square free, and hence whenever $a_{i}=a_{j}$ for $i\neq j$, one has $b_{i}\neq b_{j}$,
  • $\bullet$
    $a_{i}\neq a_{1}$ for some $i\in\{2,...,n\}$,
it follows that $G(Z,Y)$ is a non-constant square-free polynomial. Thus, we can rewrite the sum above as
$\sum_{z,y\in\mathbb{F}_{q^{mr}}}\chi_{mr}(g(z))\chi_{mr}(G(z,y))=\sum_{z\in\mathbb{F}_{q^{mr}}}t_{\mathcal{L}_{\chi_{mr}(g(Z))}}(z)t_{\mathcal{H}_{G,\chi}}(z),$
Where:
  • $\bullet$
    $\mathcal{L}_{\chi_{mr}(g(Z))}$ is the Kummer sheaf over $\overline{\mathbb{A}}_{\mathbb{F}_{q^{mr}}^{1}}$ associated with the function $\chi_{mr}(g(z))$, with singularities $\mathrm{Sing}(\mathcal{L}_{\chi_{mr}(g(Z))})=\{\infty\}\cup\{-b_{i}:i\text{ such that }a_{i}=a_{1}\}$.
  • $\bullet$
    the sheaf $\mathcal{H}_{G,\chi_{mr}}$ is the one in Example 5.3 associated to the polynomial $G(Y,Z)$. In particular, $\operatorname{sing}(\mathcal{H}_{G,\chi_{mr}})\cap\overline{\mathbb{A}}_{\mathbb{F}_{q}}^{1}(\overline{\mathbb{F}}_{q})=\{z\in\overline{\mathbb{A}}_{\mathbb{F}_{q}}^{1}(\overline{\mathbb{F}}_{q}):G(z,Y)\text{ is not square-free}\}$.
We claim that the sheaf $\mathcal{H}_{G,\chi_{mr}}$ is lisse at $b_{1}$. To do this, in is enough to show that the polynomial $G(-b_{1},Y)$ is square-free: for distinct $i,j$ such that $a_{j}\neq a_{1}$, the factors
$(a_{i}-a_{1})Y+(b_{i}-b_{1}),\quad(a_{j}-a_{1})Y+(b_{j}-b_{1})$
are proportional only if
$\frac{a_{i}-a_{1}}{a_{j}-a_{1}}\cdot(b_{j}-b_{1})\neq b_{i}-b_{1},$
but this is not possible thanks to hypothesis $(ii)$. Since $b_{1}\in\mathrm{Sing}(\mathcal{L}_{\chi_{mr}(g(Z))})$ but $b_{1}\not\in\operatorname{sing}(\mathcal{H}_{G,\chi_{mr}})$ and $\mathcal{L}_{\chi_{mr}(g(Z))}$ is irreducible, the sheaf $\mathcal{L}_{\chi_{mr}(g(Z))}$ does not appear in the decomposition $\mathcal{H}_{F,\chi}=\bigoplus_{j}\mathcal{F}_{j}$. Then, the semi-orthogonality of trace functions ([FKMS19, Theorem $5.2$]) implies that
$\sum_{x,y\in\mathbb{F}_{q^{mr}}}\chi_{mr}(F(x,y))=\sum_{z\in\mathbb{F}_{q^{mr}}}t_{\mathcal{L}_{\chi_{mr}(g(Z))}}(z)t_{\mathcal{H}_{G,\chi_{mr}}}(z)\ll_{n}q^{mr},$
as claimed. ∎
We are finally ready to prove:
Proposition $2.4$.
Let $F(X,Y)\in\mathbb{F}_{p}[X,Y]$ a square free polynomial such that
$F[X,Y]=\prod_{i=1}^{n}(X+a_{i}Y+b_{i})\text{ over }\overline{\mathbb{F}}_{p},$
and assume that the following conditions hold:
  • $i)$
    There exists $i\in\{2,...,n\}$ such that $a_{1}\neq a_{i}$.
  • ii)ii)
    $\frac{a_{i}-a_{1}}{a_{j}-a_{1}}\cdot(b_{j}-b_{1})\neq b_{i}-b_{1}$ for any $i,j$ such that $a_{j}\neq a_{1}$ and $i\neq j$.
  • $iii)$
    For every $i,j$ with $i\neq j$, $b_{i}\neq b_{j}$.
Moreover, assume that $a_{i}\neq 0$ for every i=1,,ni=1,...,n. Let $s\in\mathbb{F}_{p}$, then there exists $A_{s}\subset\{(r,t)\in\mathbb{F}_{p}\times\mathbb{F}_{p}\}$ such that
  • $i)$
    for any $(r,t)\in(\mathbb{F}_{p}\times\mathbb{F}_{p})\setminus A_{s}$ and , one has
    $\left|\sum_{(x,y,z)\in\mathbb{F}_{p}^{3}}\left(\frac{F(x,y)}{p}\right)\left(\frac{F(z,y+r)}{p}\right)e_{p}(s(x-z)+ty)\right|\ll_{n}p^{3/2}(s,t,p)^{1/2}.$
  • ii)ii)
    $A_{s}=O_{n}(1)$.
Proof.
The case $s=t=0$ is a special case of Lemma 5.7. Hence, we may assume $(s,t)\neq(0,0)$. We start by proving the case when $s\neq 0$. We begin by noticing that:
$\sum_{x,y,z\in\mathbb{F}_{p}^{3}}\left(\frac{F(x,y)}{p}\right)\left(\frac{F(z,y+r)}{p}\right)e_{p}(s(x-z)+ty)=\sum_{\begin{subarray}{c}w,x,y,z\in\mathbb{F}_{p}^{4}\\ w^{2}=F(x,y)F(z,y+r)\end{subarray}}e_{p}(s(x-z)+ty).$
Then, thanks to Theorem 5.5, it is enough to estimate
$\begin{split}N_{m}(k)&=|\{(w,x,y,z)\in\mathbb{F}_{p^{m}}:w^{2}=F(x,y)F(z,y+r),\text{ }s(x-z)+ty=k\}|\\ &=|\{(w,x,y)\in\mathbb{F}_{p^{m}}:w^{2}=F(x,z)F(x+\overline{s}(ty-k),y+r)\}|,\end{split}$
for any $m\geq 1$ and any $k\in\mathbb{F}_{p^{m}}$. First observe that over $\overline{\mathbb{F}}_{p}$ we can decompose $F(X,Y)F(X+\overline{s}(tY-k),Y+r)$ as
$\begin{split}F(X,Y)F(X+\overline{s}(tY-k),Y+r)&=\prod_{i=1}^{n}(X+a_{i}Y+b_{i})\prod_{i=1}^{n}\left(X+(a_{i}+\overline{t}s)Y+b_{i}+a_{i}r-\overline{t}k\right).\end{split}$ (25)
[…]
$X+a_{i}Y+b_{i},\quad X+(a_{\alpha(i)}+\overline{s}t)Y+b_{\alpha(i)}+a_{\alpha(i)}r-\overline{s}k,$
are proportional. From this we deduce that for every i=1,,ni=1,...,n one has
$\begin{cases}s(a_{i}-a_{\alpha(i)})-t=0\\ s(b_{i}-b_{\alpha(i)}-a_{\alpha(i)}r)=k.\end{cases}$
Now let $i=\alpha_{1}$ and $j$ such that $\alpha(j)=1$. Then $s$ and $t$ are the solutions of the system
$\begin{cases}s(a_{1}-a_{i})-t=0\\ s(b_{1}-b_{i}-a_{i}r)=k\\ s(a_{j}-a_{1})-t=0\\ s(b_{j}-b_{1}-a_{1}r)=k.\end{cases}$
Hence, from the system above, we deduce that
$\begin{cases}t=s(a_{1}-a_{i}),\\ r=\frac{2b_{1}-b_{i}-b_{j}}{a_{1}-a_{j}}\end{cases}$
Now, if we define
$A_{s}=\left\{(t,r)\in(\mathbb{F}_{p}^{\times})^{2}:t=s(a_{1}-a_{i}),\quad r=\frac{2b_{1}-b_{i}-b_{j}}{a_{1}-a_{j}}\text{ for any }i,j=1,....,n\text{ s.t. }a_{i}\neq a_{1}\right\},$
it follows that for any $(t,r)\not\in A_{s}$ and any $k\in\mathbb{F}_{p^{m}}$ the polynomial $W^{2}=F(X,Y)F(X+\overline{s}(tY-k),Y+r)$ is irreducible for any $k$. Hence, using the Lang-Weil bound, we get that
$N_{m}(k)=p^{2m}+O_{n}(p^{3m/2}),$
for any $k\in\mathbb{F}_{p^{m}}$, provided that $(t,r)\not\in A_{s}$. From now on, we always assume that $(t,r)\not\in A_{s}$. Moreover, to lighten the notation, we denote by $s^{\prime}=\overline{s}t$, $k^{\prime}=\overline{s}k$ The next step is to show that the polynomial $F(X,Y)F(X+s^{\prime}Y-k^{\prime},Y+r)$ satisfies the hypothesis of Lemma 5.8 for all but $O_{n}(1)$ choices of $k$. Notice that the polynomials $F(X,Y)$ and $F(X+s^{\prime}Y-k^{\prime},Y+r)$ both satisfy the hypothesis of Lemma 5.8: the polynomial $F(X,Y)$ by hypothesis, and $F(X+s^{\prime}Y-k^{\prime},Y+r)$ because is obtained by $F(X,Y)$ using a change of variables.
[…]
$X+a_{i}Y+b_{i},\quad X+(a_{j}+s^{\prime})Y+b_{j}+a_{j}r-k^{\prime},$ (26)
are not proportional, which is guaranteed if $k^{\prime}\neq b_{i}-b_{j}+a_{j}r$.
Next, we verify that the polynomial $F(X,Y)F(X+s^{\prime}Y-k^{\prime},Y+r)$ also satisfies the second condition in Lemma 5.8 for all but $O_{n}(1)$ many $k^{\prime}$. Since $F(X,Y)$ and $F(X+s^{\prime}Y-k^{\prime},Y+r)$ already satisfy this condition individually by hypothesis, it remains to ensure that any pair of linear factors coming from different components of the product still satisfy condition $(ii)$ of Lemma 5.8. That is, we must check that given two such factors,
$X+a_{i}Y+b_{i},\quad X+(a_{j}+s^{\prime})Y+b_{j}+a_{j}r-k^{\prime},$
with $a_{i}\neq a_{1}$ and $a_{j}+s^{\prime}\neq a_{1}$, they satisfy
$\frac{b_{j}+a_{j}r-k^{\prime}-b_{1}}{a_{j}+s^{\prime}-a_{1}}=\frac{b_{i}-b_{1}}{a_{i}-a_{1}},$ (27)
[…]
Hence, if we denote by $\chi_{m}$ the unique character of order 22 over $\mathbb{F}_{p^{m}}$, it follows from Lemma 5.8 that
$N_{m}(k)=p^{2m}+\sum_{x,y}\chi_{m}(F(x,y)F(x+s^{\prime}y-k^{\prime},y+r))=p^{2m}+O_{n}(p^{m}),$
for all but $O_{n}(1)$ values of $k$. Thus, we get the result by applying Theorem $5.$
We now deal with the case $s=0$, $t\neq 0$. In this case, the exponential sum becomes
$\sum_{x,y,z\in\mathbb{F}_{p}^{3}}\left(\frac{F(x,y)}{p}\right)\left(\frac{F(z,y+r)}{p}\right)e_{p}(ty)=\sum_{\begin{subarray}{c}w,x,y,z\in\mathbb{F}_{p}^{4}\\ w^{2}=F(x,y)F(z,y+r)\end{subarray}}e_{p}(ty).$
With the goal of using Hooley’s Theorem, for any $m\geq 1$ and any $k\in\mathbb{F}_{p^{m}}$ we have to count
$\begin{split}N_{m}(k)&=|\{(w,x,y,z)\in\mathbb{F}_{p^{m}}:w^{2}=F(x,y)F(z,y+r),\text{ }sy=k\}|\\ &=|\{(w,x,y)\in\mathbb{F}_{p^{m}}:w^{2}=F(x,\overline{t}k)F(z,\overline{t}k+r)\}|\\ &=p^{2m}+\sum_{(x,z)\in\mathbb{F}_{p^{m}}^{2}}\chi_{m}(F(x,\overline{t}k)F(z,\overline{t}k+r))\\ &=p^{2m}+\left(\sum_{x\in\mathbb{F}_{p^{m}}}\chi(F(X,\overline{t}k)\right)\left(\sum_{z\in\mathbb{F}_{p^{m}}}\chi_{m}(F(z,\overline{t}k+r))\right).\end{split}$
We claim that for every value of $k\neq 0$ the polynomials $F(X,\overline{t}k)$ and $F(Z,\overline{t}k+r)$ are not squares. Let us first show how from the claim we can conclude the proof of the proposition: Since $F(X,\overline{t}k)$ and $F(Z,\overline{s}k+r)$ are not square for every $k$, we have ([IK04, Theorem $11.23$])
$\left|\sum_{x\in\mathbb{F}_{p^{m}}}\chi(F(X,\overline{t}k)\right|\ll_{n}p^{\frac{m}{2}},\quad\left|\sum_{z\in\mathbb{F}_{p^{m}}}\chi_{m}(F(z,\overline{t}k+r))\right|\ll_{n}p^{\frac{m}{2}}$
Hence,
$N_{m}(k)=p^{2m}+O_{n}(p^{m}),$
for every $k$. Hence, applying Theorem 5.5, we get
$\left|\sum_{x,y,z\in\mathbb{F}_{p}^{3}}\left(\frac{F(x,y)}{p}\right)\left(\frac{F(z,y+r)}{p}\right)e_{p}(ty)\right|\ll_{n}p.$
It rests to prove the claim. To do this it is enough to show that $F(X,\overline{t}k)$ is not a square for every value of $k$. Since
$F(X,\overline{t}k)=\prod_{i=1}^{n}(X+\overline{t}ka_{i}+b_{i})\text{ over }\overline{\mathbb{F}}_{p}$
the polynomial $F(X,\overline{t}k)$ is a square if and only if there exists $\alpha\in S_{n}\setminus\{\operatorname{Id}\}$ such that for every i,,ni,...,n the factors
$X+\overline{t}ka_{i}+b_{i},\quad X+\overline{t}ka_{\alpha(i)}+b_{\alpha(i)}\text{ are equal for every }i\in\{1,...,n\},$
that is if and only if
$k(a_{i}-a_{\alpha(i)})=t(b_{\alpha(i)}-b_{i})\text{ for every }i\in\{1,...,n\},$ (28)
Now we distinguish two cases:
  1. $a)$
[…]
    $k=t\frac{b_{i}-b_{1}}{a_{i}-a_{1}}=t\frac{b_{j}-b_{1}}{a_{j}-a_{1}},$
    which would be in contradiction with assumption $(ii)$.
  2. $b)$
[…]
    but this is not possible since $\alpha\neq\operatorname{Id}$ and our assumption $(iii)$.
Acknowledgments
This project began while the author was affiliated with University of Basel, and it originated as a joint project with Pierre Le Boudec. Although P. Le Boudec has decided to moved away from academia and preferred not to be listed as a co-author, the author is deeply grateful for his contribution to the paper. The author would also like to thank Julian Demeio for helpful discussions, which led to a simplification of Lemma 4.3. Thanks are due as well to Tim Browning and Lillian Pierce for their valuable comments on an earlier version of the paper, which helped improve the exposition of the material.